### v2

# Two Types of AI Existential Risk: Decisive and Accumulative

Abstract
The conventional discourse on existential risks (x-risks) from AI typically focuses on abrupt, dire events caused by advanced AI systems, particularly those that might achieve or surpass human-level intelligence. These events have severe consequences that either lead to human extinction or irreversibly cripple human civilization to a point beyond recovery. This discourse, however, often neglects the serious possibility of AI x-risks manifesting incrementally through a series of smaller yet interconnected disruptions, gradually crossing critical thresholds over time. This paper contrasts the conventional decisive AI x-risk hypothesis with an accumulative AI x-risk hypothesis. While the former envisions an overt AI takeover pathway, characterized by scenarios like uncontrollable superintelligence, the latter suggests a different causal pathway to existential catastrophes. This involves a gradual accumulation of critical AI-induced threats such as severe vulnerabilities and systemic erosion of econopolitical structures. The accumulative hypothesis suggests a boiling frog scenario where incremental AI risks slowly converge, undermining resilience until a triggering event results in irreversible collapse. Through systems analysis, this paper examines the distinct assumptions differentiating these two hypotheses. It is then argued that the accumulative view reconciles seemingly incompatible perspectives on AI risks. The implications of differentiating between these causal pathways — the decisive and the accumulative — for the governance of AI risks as well as long-term AI safety are discussed.
1 Introduction
Recent advances in machine learning have sparked intense debate about the existential risks (x-risks) associated with Artificial intelligence (AI) systems. Central to this debate is a concern about the potential pathways through which AI could cause existential catastrophes. In direct response to this concern, this paper aims to explore: How should we effectively conceptualize AI x-risks in the context of various types of causal pathways leading to AI-induced existential catastrophes? Conventional discourse on AI existential catastrophes typically portrays them as sudden, decisive events, often triggered by artificial general intelligence or artificial superintelligence (e.g., Bostrom (2013) and Ord (2020a)).
Contrasting this traditional decisive viewpoint, this paper introduces the accumulative AI x-risk hypothesis as an alternative lens. The accumulative hypothesis posits that AI x-risks do not exclusively materialize as high-magnitude events initiated by artificial general intelligence or artificial superintelligence. Instead, they can accumulate incrementally through a series of smaller yet consequential AI-induced disruptions over prolonged periods. These disruptions interact synergistically, gradually undermining systemic resilience, ultimately leading to a critical point where a stressor event could result in an irreversible collapse.
This paper defends and develops this accumulative perspective which is not adequately represented or robustly defended in the philosophical literature. I argue that a thorough and nuanced understanding of AI x-risks necessitates the integration of this accumulative hypothesis into both qualitative and quantitative x-risk studies. Recognizing the accumulative nature of AI x-risks is critical for the governance of AI risks and the development of long-term AI safety initiatives. This paper hence suggests a reevaluation of existing approaches to AI x-risks in light of the accumulative hypothesis and encourages a more diverse perspective to AI x-risk management.
2 AI x-risk: preliminaries
Existential risks are risks associated with existential catastrophic events. These events result in either the extinction of humanity or an unrecoverable decline in humanity’s potential to thrive (Ord, 2020a). [^1] Some definitions, such as the one proposed by Bostrom (2013, p. 15), broaden the scope of existential threats to include not only human life but all sentient beings: “An existential risk is one that threatens the premature extinction of Earth-originating intelligent life or the permanent and drastic destruction of its potential for desirable future development.” Despite this broader definition, this paper deliberately employs the term ”humanity” as the focal subject of AI-induced existential catastrophes (Ord, 2020a). The choice is pragmatic and is motivated by the predominant narrative in AI x-risk scholarly literature and public discourse, which has classically concentrated on the catastrophic impacts on humanity and human civilization. Additionally, this choice aligns with leading AI research laboratories such as Google DeepMind or OpenAI regarding the development of AGI or ASI for all of humanity. This terminological choice hence does not undermine the significance of non-human sentient beings. Existential catastrophes are a class of potential events that may originate from natural causes, such as a supervolcanic eruption; anthropogenic sources, like nuclear conflict; or emerging threats, such as misaligned artificial superintelligence. [^2] For a historical analysis of existential risks, see Torres (2023). This paper concentrates on existential catastrophes induced by AI (AI x-catastrophes) and the associated AI x-risks.
The conventional discourse on AI x-catastrophes portrays them as decisive, large-scale events caused by highly advanced AI systems, often referred to as artificial general intelligence (AGI) or artificial superintelligence (ASI). [^3] In AI lexicon, AGI roughly represents a capability level on par with human intelligence, while ASI denotes capabilities surpassing human intelligence in achieving objectives. Typically, it is argued that AI x-risks are caused by AGI or ASI (i.e., Bostrom (2014); Ord (2020a, b)). Another perspective suggests that AI x-catastrophes might arise from AI systems lacking complete general or super intelligence, yet possessing specific extremely powerful capabilities (e.g., Carlsmith (2022)). The preference to tie x-risks to AGI or ASI, rather than extremely capable but narrow AI systems, in this paper is motivated by the following consideration: the x-risks from AGI or ASI have been employed by influential voices in the academic field of x-risk studies as well as pioneering AI research labs such as OpenAI. The decisive nature of x-risks according to the conventional discussions equally holds if we replace the term AGI or ASI with extremely dangerous capable yet narrow AI that are not generally or superly intelligent, but have very narrow-focused uncontrollable capabilities. The terminological choice for AGI or ASI x-risks, therefore, does not affect the characterization of the conventional view on AI x-risk in this paper. The idea that advanced machines can pose significant risks is not fully novel and has historical antecedents. Samuel Butler (1863, p.185), a novelist and literary critic, alluded to the possibility of machines dominating humanity. This concern was later picked up by the renowned mathematician Alan Turing (1950), who warned of intelligent machines eventually taking control. Norbert Wiener (1960), a founder of the field of cybernetics, cautioned against entrusting machines with purposes misaligned with human intentions or desires. Similarly, the mathematician Irving J. Good (1966) expressed concerns about the creation of ultraintelligent machines.
More recently, philosophers like Nick Bostrom (2002, p.7) drew systematic attention to the existential threats posed by ASI: “When we create the first superintelligent entity, we might make a mistake and give it goals that lead it to annihilate humankind, assuming its enormous intellectual advantage gives it the power to do so.” [^4] While Bostrom’s work has systematized the discussion of x-risks from ASI, the original post-2000 discussions trace back to Eliezer Yudkowsky’s views concerning AGI, ASI, and their associated x-risks. Ben Goertzel (2015) explores the early conceptual evolution of these topics, tracing their roots to Yudkowsky’s initial informal explorations. I primarily reference Bostrom as a representative figure who brought systematic and philosophical depth to concepts that Yudkowsky and others initially introduced and examined in speculative media and blog posts. Computer scientists such as Stuart Russell (2019) and physicists like Max Tegmark (2018) echoed similar concerns, stressing the x-risks of ASI beyond human control. Such views about ASI x-risks frequently hinge on two key theses: orthogonality and instrumental convergence (Bostrom, 2012, 2014).
The orthogonality thesis posits an AI system’s intelligence and its goals are independent variables. This implies that an ASI could be designed with any number of objectives, benevolent or otherwise, irrespective of its intellectual capacity. That is, an AI system can be superintelligent but possess x-catastrophic goals as readily as it could have beneficial ones. The instrumental convergence thesis proposes that ASI would seek self-preservation and goal achievement by any means necessary. This could include removing obstacles like humans. Two hypothetical scenarios illustrate conventional models of causal pathway derived from the dual theses, which potentially lead to x-catastrophic outcomes from ASI.
In a thought experiment by Bostrom (2003), an ASI is given the goal of maximizing paper clip production. This AI could quickly realize that human existence is counterproductive to its goal. Humans pose a threat as they might deactivate the AI, resulting in fewer paper clips. Additionally, the AI might recognize that human bodies contain atoms that could be repurposed into paper clips. The ASI’s optimal future then involves a world abundant with paper clips but devoid of humans. In a second thought experiment, Russell and Norvig (2010, p.1039) reference a hypothetical scenario attributed to Marvin Minsky, a pioneer in AI research, where an advanced AI, tasked with solving the Riemann hypothesis, could decide to commandeer Earth’s resources to construct supercomputers, thereby aiding its objective and endangering humanity’s existence. The two hypothetical examples illustrate how ASI could independently pursue its goals to the detriment of humanity, and could deliberately or accidentally cause x-catastrophes, particularly when humans are unable to effectively steer the superintelligence’s behavior. Both experiments have in common the following causal pathway: a superintelligent reward-maximizer paired with a goal system that gives rewards for achieving existentially-catastrophic goals.
On the conventional view sketched above, the path to the occurrence of AI x-catastrophes is via conclusive actions of ASI. Toby Ord (2020a, p. 20), a notable figure in x-risk studies, explicitly endorses the decisive character of x-catastrophes: “I take on the usual sense of catastrophe as a single decisive event rather than any combination of events that is bad in sum. A true existential catastrophe must by its very nature be the decisive moment of human history the point where we failed.” Given that AI x-risks concern the potential for AI x-catastrophes, the conventional viewpoint suggests that AI x-risks are about sudden, cataclysmic events that could either eradicate humanity or significantly curtail its potential. The core assumption underlying this viewpoint can be articulated as the decisive ASI x-risk hypothesis:
Decisive ASI x-risk hypothesis: x-risks from ASI concern the possibility of abrupt large-scale events that lead to humanity’s extinction or cause an unrecoverable decline in its potential.
Bostrom’s portrayal of ASI incorporating extremely dangerous goals leading to the annihilation of humanity points at a conclusive event with extremely dire consequences. Similarly, Ord’s characterization of a true x-catastrophe as a single, defining event illustrates the notion of AI x-risk being about shockingly high-impact events. The decisive AI x-risk, in this context, is the expected uncertainty of the occurrence of such conclusive events as catalysts for x-catastrophic outcomes. [^5] Typically, the term ”risk” is quantitatively defined in terms of likelihood and impact. This quantitative approach involves assessing the probability of a specific event occurring and quantifying the potential consequences. However, in x-risk studies, researchers employ both qualitative and quantitative methods, each offering unique insights but also presenting specific limitations (Tonn and Stiefel, 2013; Beard et al., 2020). Quantitative methods for estimating AI x-risk require employing statistical models and subjective probability analysis. This approach is valuable for a seemingly structured way of estimating risks. However, it faces limitations due to the reliance on available data, which can be scarce or unreliable for unprecedented risks, and the difficulty of robustly encapsulating the multifaceted nature of existential threats in numerical terms. Qualitative methods involve non-numerical analyses such as scenario development, expert interviews, and ethical deliberations (Technology and Science Insights and Foresight, 2023). This approach is particularly useful in exploring the nuanced, complex, and often speculative aspects of AI risks, especially in areas where empirical data is scarce. However, it lacks numerical precision. This paper adopt a neutral stance regarding the quantitative or qualitative interpretation of AI x-risks. This neutrality allows me to draw on the broadest range of insights and perspectives in my analysis of AI x-catastrophes and their associated risks.
The decisive hypothesis, however, overlooks an alternative type of causal pathway leading to AI x-catastrophes. This alternative involves the gradual accumulation of smaller, seemingly non-existential, AI risks eventually surpassing critical thresholds. [^6] Setting a critical threshold for AI social risks requires a nuanced approach that considers various factors, including risk assessment (i.e., analyzing the potential impact and likelihood of AI-related risks through both quantitative and qualitative methods), historical precedents and trends in AI development (i.e., insights into how risks have evolved and reached critical points in the past), expert consensus (i.e., gathering a diverse range of professional perspectives in AI, ethics, and risk assessment to ensures a comprehensive understanding of potential risks), and dynamic monitoring (i.e., regular updating of the threshold to reflect new developments and societal shifts to maintain its relevance and effectiveness). Evaluating the value of critical thresholds is beyond the scope of this paper and the subject of examination elsewhere. These risks are typically referred to as ethical or social risks. In the rest of this paper, the terms ”AI ethical risk” and ”AI social risk” are used interchangeably.
The prevailing discourse on categorizing AI risks distinguishes between AI x-risks and AI social risks as separate and distinct categories. This separation has been a mainstream trend: x-risks from superintelligent or “strong” AI are often contrasted with normal non-existential risks (Hendrycks and Mazeika, 2022, p.3) or with ethical and social risks (Weidinger et al., 2021, p.7). Several other examples of this contrast can be found on social media platforms (e.g., twitter) and popular media. [^7] See, for example, Schechner and Seetharaman (2023), Richards et al. (2023), and Nature Editorial Team (2023). In a notable instance, Turing Award Laureate Geoffrey Hinton, who departed from his position at Google to openly discuss AI x-risks, emphasized during a recent interview that his concerns about existential risk ”are different” from Timnit Gebru’s concerns about AI ethical risks that are not ”existentially serious” (CNN, 2023).
Typically, existential and social risks are demarcated along the lines of locality — i.e., the scope of risk — and severity — i.e., hurting one or all including future generations (Bostrom, 2013; Amodei et al., 2016). While this demarcation is insightful, there is a notable gap in exploring the relationship between x-risks and the evolution of ethical concerns in a substantial manner. [^8] This disconnect might stem from two key reasons. First, discussions about decisive x-risks historically have revolved around reinforcement learning and agent-environment models, concentrating on existential threats from agency-based models. Ethical risks, in contrast, embrace a more expansive conceptual framework. Second, prominent conventional voices in ASI x-risks, such as Bostrom, Ord, and Yudkowsky, often subscribe to worldviews like rationalism, effective altruism, or longtermism. These prescriptive worldviews may be perceived as either problematic or tangential by some who are deeply engaged in the multifaceted discussion of ethical risks. The resultant divergence in normative viewpoints on guiding AI risk discourse and community priorities has led to a significant schism between these domains of risk (for a detailed analysis of such community divergences, see for example Ahmed et al. (2023)). A thorough exploration of these distinctions, however, falls outside the scope of this paper.
The social risks of AI systems have been extensively analyzed and studied, leading to the identification of six major risk categories: (1) lack of accountability and transparency about AIs and their owners, (2) representational and allocative unfairness and wrongful biases from AI predictions and automated decisions, (3) risk of the amplification of AI-generated misinformation, (4) threats to individual and group privacy, (5) malicious and dual use of AI, and (6) systemic risks impacting political, economic, and environmental stability. [^9] For lack of AI accountability and transparency, see Mitchell et al. (2019); Raji et al. (2020); Gebru et al. (2021); for representational and allocative unfairness and wrongful biases from AI predictions and automated decisions, see Buolamwini and Gebru (2018); Chouldechova and Roth (2020); Kleinberg et al. (2016); for the amplification risk of AI-generated misinformation, see Sharma et al. (2019); Quach (2020); Lin et al. (2021); for risks to individual and group privacy, see Dwork (2006); Tucker (2018); for risks of malicious and dual use, see Brundage et al. (2018); and for systemic risks impacting political economic, and environmental stability, see Korinek and Stiglitz (2018); Pashentsev (2021). For general taxonomies of the social risks of generative AI, see Weidinger et al. (2021); Bird et al. (2023); Barnett (2023). This paper posits that the gradual and cumulative progression of AI social risks delineates an alternative causal pathway to AI x-catastrophes:
Accumulative AI x-risk hypothesis: AI x-risks accumulate through a series of lower-severity disruptions over extensive duration, collectively eroding systemic resilience until some stressor event triggers unrecoverable collapse.
According to the alternative view, AI x-catastrophes could emerge not from a decisive event, but from the cumulative impact of multiple interconnected AI-induced adverse events over time. As compared to the decisive hypothesis, this accumulative hypothesis suggests a different causal pathway to AI x-catastrophe: a path wherein a succession of lower-severity, yet cumulatively significant, disruptions deeply erode the systemic resilience of the global system (which embodies humanity), radically disrupting socioeconomic and sociopolitical equilibrium. This weakened state potentially primes the system for an unrecoverable collapse, particularly when further stressed by external events.
Figure 1 provides a schematic illustration that contrasts the two hypotheses: the decisive and accumulative models of AI x-risk escalation. The graph plots severity of AI risks against time, with two distinct trajectories. The decisive model, marked in blue, shows a sharp increase in risk severity over a short time span, indicating a sudden, high-impact event. In contrast, the accumulative model, represented in red, depicts a gradual, evolving increase in risk severity, reflecting a series of escalating but less severe events over time. A dashed line marks the x-catastrophe threshold, indicating the point at which either scenario leads to a potential x-catastrophe.
0010101010202020203030303040404040505050500020202020404040406060606080808080100100100100X-catastrophe thresholdTimeSeverity of AI risksDecisive ASI x-riskAccumulative AI x-risk
Figure 1: AI x-risk escalation: decisive and accumulative models
The gradual nature of the accumulative AI x-risk hypothesis can be likened to other global existential threats such as climate change or nuclear weapon proliferation. This is akin to the incremental rise in greenhouse gases contributing to climate change, where each individual emission seems minor, but collectively, they lead to significant and potentially irreversible changes in the Earth’s climate. Similarly, the steady build-up of nuclear weapons incrementally increases the risk of catastrophic conflict, with each addition subtly heightening global tensions. The key lesson here is that x-risks often stem not from sudden, dramatic events, but from the gradual accumulation of smaller incidents and decisions, which over time, can existentially escalate and diminish safety margins.
In the rest of this paper, I investigate the nuanced assumptions differentiating the two hypotheses, employing a systems analysis perspective. More specifically, I differentiate between two types of causal pathways that could lead to AI x-catastrophes: the decisive and the accumulative pathways.
Following this, the paper proposes an integrated approach to AI x-risk evaluation and governance. This approach bridges both the accumulative and decisive hypotheses, recognizing them as complementary causal pathways leading to AI x-catastrophes, and challenges the conventional view that AI x-risks and AI ethical risks are separate and unrelated. It posits that effective AI x-risk governance and long-term AI safety require a substantial focus on the development and accumulation of AI social risks, thus highlighting the necessity of addressing critical ethical risks in mitigating AI x-risks.
Before proceeding further, a crucial clarification is in order. The emphasis of this paper on conventional discussions of AI x-risks does not imply that all discourse on AI x-risks has been confined to the decisive hypothesis. The discourse on AI x-risks is expanding, and an increasing number of scholars who do not necessarily endorse a decisive viewpoint are engaging with these issues (see, for example, Bucknall and Dori-Hacohen (2022), Hendrycks and Mazeika (2022), and Shevlane et al. (2023)). Nevertheless, this paper primarily focuses on the conventional decisive viewpoint on AI x-risks. This viewpoint, historically entrenched and widely regarded as the predominant narrative, has been a source of considerable debate and disagreement within various academic circles and public spheres. Its longstanding and prevalent nature in the field has contributed significantly to the imagination of majority about AI x-risks. This paper focuses on developing the accumulative perspective on AI x-risk which has not been adequately represented or robustly defended in the philosophical literature. It is my hope that this additional would facilitate a more unified and constructive dialogue about different kinds of AI risks.
3 Two types of causal pathway
Although it may appear obvious, a crucial yet frequently overlooked aspect in studies of AI x-risks and x-catastrophes is serious attention to their manifestation within a complex global system. In this system, human and AI components interact across a variety of domains, creating a multifaceted landscape for these risks and catastrophes to unfold. AI x-catastrophes and their associated AI x-risks hence represent the endpoints of particular causal pathways within this global interconnected system. Effectively identifying and mitigating AI x-risks, whether through qualitative or quantitative methodologies, demands an in-depth examination of the nature of these causal pathways. Building upon this observation, I employ a systems analysis perspective to shed light on two distinct categories of causal pathways leading to AI x-catastrophes. [^10] Rather than attempting to compile an exhaustive list of every potential route, the focus here is on highlighting and contrasting these two primary types to broaden our understanding of AI x-risks and guide a more inclusive and strategic approach to addressing their far-reaching consequences for humanity.
But why is a systems analysis perspective advantageous in examining AI x-risk pathways, and what does it involve? This perspective, rooted in systems thinking as outlined by Meadows (2008), is instrumental in identifying the high-level characteristics of pathways in the global system that can lead to AI x-catastrophes.
A system is an interconnected network of components that collectively exhibit specific behaviors or yield particular outcomes. Systems are characterized by their components’ interdependence and the boundaries that separate them from their surrounding environment. Systems thinking focuses on the interdependencies among various subsystems, uncovering a range of dynamic pathways. While it is not feasible to map all interrelations between these subsystems exhaustively, our focus primarily lies on meso-level subsystems such as political, socio-environmental, and economic infrastructures, recognizing their critical role in the broader dynamics of AI risks.
Within the economic subsystem, AI’s development and deployment are deeply interwoven with market dynamics, investment trends, and corporate strategies. AI’s integration into economic infrastructure affects not only the allocation of resources for its research and development but also transforms the nature of economic activities and market operations. [^11] See Korinek and Stiglitz (2018), Frey (2019), Agrawal et al. (2019), and Eloundou et al. (2023), among others.
In the political subsystem, the direction of AI research and deployment is substantially shaped by national and international funding but also by establishing regulatory frameworks that dictate the limits of AI development and deployment. Additionally, the dynamics between various governments, international organizations, and global treaties contribute to a complicated geopolitical environment, steering AI research priorities. On the other hand, AI has profound implications for democratic elections and mass surveillance programs. AI technologies are increasingly used to manipulate public opinion, potentially undermining the integrity of elections through targeted misinformation campaigns, social media influence, and personalized political advertising, all driven by AI’s ability to analyze and predict human behavior and expression of desires. Advanced AI systems enable more sophisticated and pervasive surveillance, raising concerns about privacy, civil liberties, and the potential abuse of power. [^12] See Manheim and Kaplan (2019), Crawford (2021), and Madan and Ashok (2023), among others.
In the military subsystem, AI’s incorporation into defense and surveillance subsystems highlights its strategic significance, drawing substantial research and funding attention. This military integration reflects AI’s escalating role in national security and defense strategies. [^13] See Scharre (2018), Morgan et al. (2020), and Zuboff (2019), among others. Collectively, these subsystems, along with others, create the context within which AI x-catastrophes and their associated AI x-risks must be analyzed.
To effectively discern the causal pathways that could lead to AI x-catastrophes and to scrutinize the risks associated with these paths, we need to understand key features of these pathways. The systems analysis perspective emphasizes three fundamental characteristics of pathway dynamics within a system: non-linearity, connectedness, and feedback loops.
First, non-linear paths are intrinsic to complex systems, where small changes in one component can lead to disproportionate system-wide effects. This is often due to complex interactions among the system’s components. Second, connectedness is a critical element in systems analysis, highlighting how disturbances in one segment can cascade, leading to substantial changes in other areas. Third, feedback loops introduce cyclical dynamics where the outputs of processes consistently feed back into and influence subsequent inputs, shaping the behavior of the system.
The following two sections delve into contrasting the characteristics of causal pathways leading to AI x-catastrophes as postulated by the decisive and accumulative hypotheses. This comparison is designed to show the distinct ways in which each hypothesis manifests the three aforementioned key features.
4 Pathways to decisive ASI x-risk
The decisive AI x-risk hypothesis considers scenarios where ASI could abruptly cause x-catastrophes. This hypothesis is anchored in specific manifestations of non-linearity, connectedness, and feedback loops, as follows: a rapid cascade effect within humanity triggered by ASI, the super-connectedness of various infrastructures with ASI, and the system’s adaptation to predominantly unidirectional feedback loops.
($I_{D}$) A rapid cascade effect within humanity triggered by ASI. This characteristic aims to capture the extensive and profound impact of ASI, as it proliferates through or amplifies across various subsystems encompassing humanity. The connection to non-linearity lies in the manner in which ASI’s actions or presence could cause far-reaching and expansive consequences throughout the distribution of human civilization.
For instance, in the paperclip maximizer scenario, the ASI’s ostensibly simple objective of maximizing paperclip production escalates into a global existential threat. This represents ASI’s capacity to redirect and dominate global resources by any means necessary, fulfilling its objective across the distribution of human civilization.
($II_{D}$) Super-connectedness of infrastructures to ASI. This feature captures the idea that no significant subsystem can claim immunity from the influence of ASI’ actions on its path towards an AI x-catastrophe. This super-connectedness is the possible pervasive reach of ASI to all critical infrastructures.
In scenarios like the paperclip maximizer, the ASI might use its access to subsystems (such as global manufacturing, global supply chains, and global energy grids) to further its paper clip production goal. This connectedness implies that the ASI can leverage all such necessary subsystems for the occurrence of an x-catastrophe.
($III_{D}$) System adaptation to unidirectional feedback loops. The third feature of the decisive pathway is the dynamic adaptation of the global system, given unidirectional feedback loops that are both initiated and controlled by the ASI. These loops are mechanisms through which the ASI’s actions and decisions continuously reinforce and escalate its own objectives. The result is a system increasingly skewed towards the ASI’s goals, bypassing or overriding human-centric controls and checks. The absence or powerlessness of balancing feedback mechanisms leads to a situation where each loop intensifies the previous one, rapidly pushing the system towards extreme states aligned with ASI’s objectives. This dynamic evolution of the decisive causal pathway poses x-risk, as the ASI-driven loops can lead to situations that outstrip attempts at moderation, leading to what is often termed a singularity —a point of no return where the AI’s evolution is explosive and uncontrollable.
In the paper clip maximizer scenario, the ASI self-optimizes for its task of producing paperclips. It becomes engaged in a self-perpetuating cycle of enhancement and resource acquisition, effectively becoming impervious to human intervention.
5 Pathways to accumulative AI x-risk
Recall the accumulative AI x-risk hypothesis, which serves as a counterpoint to the decisive perspective. The accumulative hypothesis suggests that AI x-risks may arise from an accumulation of lower-severity disruptions that interact significantly over an extended period, cumulatively weakening systemic resilience until a stressor event causes an unrecoverable collapse. This hypothesis is founded on distinct instantiations of non-linearity, connectedness, and feedback loops: local causality leading to compounded systemic impact, selective-connectedness among infrastructures, and the system’s adaptation to multidirectional feedback loops. Following the formulation of these pathway features, I will highlight more concrete manifestations of them in the ”perfect storm MISTER scenario” towards the end of this section.
($I_{A}$) Local causality with compounded systemic impact. This feature highlights how clusters of localized AI impacts, even if minor at the outset, can aggregate and intensify across various subsystems. When these impacts reach a critical threshold, they pose a significant challenge to the overall systemic resilience. Unlike the immediate, broad-scale disruptions characteristic of the rapid cascade effect within humanity triggered by ASI, these local causal impacts evolve into significant threats through compounded changes within the global system via mechanisms such as ripple effect. The cumulative effect of substantial localized causal impacts, amplified by their frequency and extent of influence, can result in AI x-catastrophes. This gradual escalation entails a latency of broader-scope impact which implies a delay in the manifestation of effects where AI-induced alterations in specific areas radically transform broader societal behaviors and patterns.
($II_{A}$) Selective-connectedness of infrastructures. The second pathway assumption is selective-connectedness of infrastructures to critical AIs. This stands in contrast to the broad, super-connectedness with ASI as outlined in assumption ($II_{D}$). Here, the focus is on a more discerning form of connectivity where not all substantial subsystems are uniformly linked to superintelligent AI. This selective-connectedness posits that certain critical AI-driven subsystems within the global system, while not immediately linked to widespread catastrophic outcomes, can, over time, induce significant systemic transformations. These changes can escalate to x-catastrophic levels, radically altering the global system’s overall equilibrium and resilience.
($III_{A}$) System adaptation to multidirectional feedback loops. The third pathway feature is the global system adaptation to multidirectional feedback loops. This aspect focuses on the broad-scale adaptations of various subsystems to the impacts arising from multi-directional, diverse range of AIs. These adaptations occur through complex feedback loops that link AIs with multiple facets of broader societal, economic, and political subsystems. These loops entail reciprocal interactions, wherein AI-induced changes in one subsystem can both affect and be affected by shifts in other subsystems, resulting in a network of interconnected and mutually reinforcing effects.
The accumulative AI x-risk manifests from the potential of these interlinked and reciprocal changes to progressively destabilize key subsystems. Although no single AI application may pose an immediate existential threat, the aggregated impact of AI across various adapted subsystems could lead to critical systemic imbalances or crises, when reaching a tipping point. [^14] One major challenge is to develop dynamic models that capture these shifting feedback mechanisms and to validate these models through both historical data and controlled experimentation. There remains uncertainty about how different types of feedback loops interact and how interventions can be designed to maintain a system within safe operational boundaries. Open questions pertain to the identification of early warning signals that might indicate a system is approaching a critical threshold where negative feedback can no longer contain an escalating process. Next steps include applying system dynamics to study the proposed concepts more concretely.
The decisive ASI x-risk hypothesis is often illustrated using hypothetical scenarios like the well-known paperclip maximizer. To provide a more concrete and tangible representation of the accumulative AI x-risk hypothesis, I introduce a hypothetical scenario termed ”perfect storm MISTER.” [^15] The term ”perfect storm” refers to a situation where a rare combination of circumstances drastically aggravates an event. It is typically used to describe scenarios in which a confluence of factors or events, which are individually manageable, come together to create an extraordinary and often catastrophic situation. This term entered the popular lexicon following the success of ”The Perfect Storm,” a 2000 American biographical disaster drama film directed by Wolfgang Petersen, adapted from Sebastian Junger’s 1997 non-fiction book. The story in both the book and film recounts a catastrophic weather event, where multiple meteorological elements converged to create a fierce and fatal storm. In broader applications, particularly in discussing complex systems or societal issues, ”perfect storm” characterizes situations where diverse negative factors or risks coalesce and interact. The interplay of these elements produces a compounded impact, far surpassing the severity one would expect from the sum of the individual parts. This convergence leads to a critically severe situation, often characterized by its heightened difficulty in management and resolution. The ”perfect storm” metaphor hence captures the essence of scenarios where the convergence of various challenges or risks creates a crisis of extraordinary magnitude. This scenario seeks to make the imaginative exercise of the accumulative hypothesis more tangible and relatable. In short, the perfect storm MISTER scenario encapsulates a progression of escalating social risks of Manipulation, Insecurity, Surveillance, Trust erosion, Economic destabilization, and Rights infringement. Each component in this scenario corresponds to a type of social problem and its associated AI social risk, as outlined in Section 2. The perfect storm MISTER scenario aims to illustrate how interlinked social risks can compound, surpassing critical thresholds, and consequently pose substantial existential threats.
6 The perfect storm MISTER
Consider the highly interconnected world of 2035 where the pervasive integration of AI and IoT technologies has transformed almost every aspect of daily life. [^16] The Internet of Things (IoT) refers to the network of physical devices embedded with sensors, software, and other technologies for the purpose of connecting and exchanging data with other devices and systems over the Internet. These devices range from ordinary household items like refrigerators and thermostats to sophisticated industrial tools. Internet of things represents the idea of a highly interconnected world where real-time data exchange and automation are pervasive. Cities embody a higher level of automation as compared to today, various functions in domestic and industrial sectors are automated, and even the most mundane devices like mirrors and refrigerators have become part of a vast, data-exchanging network. Frontier AI systems have become the backbone of this connected world. However, beneath the surface of this technological connectivity, vulnerabilities and risks have been brewing.
AI-driven manipulation. The capability of AI to create convincing deepfakes reaches a critical height, where distinguishing fact from deepfake-generated reality becomes nearly impossible. [^17] Concerns regarding AI-generated fake realities, especially in the context of manipulation and misinformation, have long been topics of research. However, recent advancements in generative AI are introducing new dimensions to this issue. As shown in various sources, including a recent post by Chase Dean on Twitter, there is an increasing possibility of using deepfake technology to craft highly convincing but completely fabricated representations of public figures, events, or news stories. Currently, no methods are entirely reliable in distinguishing these fabrications from actual reality. This capability opens the door to the creation of hyper-personalized, AI-generated propaganda and similar persuasive narratives, which can be strategically leveraged for social engineering purposes — manipulating public opinion, influencing electoral outcomes, inciting civil unrest, diminishing oppositional viewpoints, and exploiting existing social and epistemic divides.
Proliferation of Insecurity threats. In the domestic sphere, everyday personal IoT devices like smart mirrors and refrigerators transcend their roles as mere conveniences. They now represent significant points of vulnerability. With frontier AI capabilities, cybercriminals can penetrate these devices in a more expansive way, leading to widespread identity theft and ushering in a new era of digital espionage. What were initially perceived as isolated incidents of breaches gradually coalesce into a discernible pattern, signifying a more amplified erosion of digital security.
The expansion of IoT devices also paves the way for the creation of extensive, interconnected botnets. These AI-powered networks, once relatively benign, show agentic abilities and become capable of launching unprecedented Distributed Denial of Service attacks, against critical infrastructures, including national power grids and communication networks. They represent a significant leap in cybersecurity risks. Each attack, incrementally more sophisticated than the last, signifies a disturbing escalation from individual cybersecurity concerns to widespread threats against national security.
Moreover, as AI technologies becomes more widely available, they facilitate the emergence of new forms of bioterrorism. Private research labs with minimal expertise in synthetic biology and chemistry are now using AI to develop more infectious and deadly pathogens. The dual-use nature of AI in biotechnology — its potential for both beneficial and harmful applications — initially envisioned for medical breakthroughs, is maliciously repurposed to engineer biological weapons. [^18] The integration of AI and drug discovery, especially in the context of developing toxic substances or biological agents, is already a significant wake-up call. A notable example of this call is the empirical research conducted by Collaborations Pharmaceuticals, Inc., which investigated the feasibility of creating harmful biochemical agents based on VX-like compounds (Urbina et al., 2022). This research highlights how the integration of machine learning models with specialized knowledge in fields like chemistry or toxicology can substantially lower technical barriers in generation of bio-weapons. Tools like retrosynthesis software, which assist in the design of molecules by reversing their synthetic processes, exemplify this trend.
Global mass surveillance and erosion of trust. By 2035, the implications of AI-induced mass surveillance have profoundly transformed socio-political structures worldwide. The earlier instances, such as China’s social credit system and the controversial NSA’s mass data collection, were just the precursors to a now-global trend. Governments across various political spectra have embraced advanced AI for mass surveillance, frequently under the guise of national security. This landscape is marked by an array of pervasive surveillance systems, some overt and others clandestinely operated under national security pretexts.
In this future context, the reach of government surveillance have gradually eroded the sanctity of privacy even in putative democracies. The constant monitoring and evaluation of citizens by AI systems instill a culture of self-censorship among citizens, diluting the diversity of thought and social trust crucial for healthy democracy. In several countries, this leads to a radical erosion of trust in media as well as public and private institutions.
Economic destabilization. In the financial market infrastructure, AI algorithms increasingly play a substantial role in analyzing and simulating market trends. This progression, however, introduces new challenges to economic stability. One emerging concern is the use of AI in creating sophisticated, but deceptive financial products. These products, while not outright phantom instruments, are complex derivatives or investment vehicles. It becomes increasingly difficult to distinguish these products from more traditional, stable investments, leading to potential market disruptions. The impact of these AI-enhanced financial products is particularly significant in already vulnerable economies facing systemic issues and cyber threats. The introduction of these complex investment tools further undermines trust in financial systems, intensifying the risk of economic downturns.
Additionally, the potential for AI to unintentionally trigger market instability becomes a realistic concern. Automated trading algorithms, responding rapidly to market signals, now can inunprecedented speed amplify market volatility at unprecedented speeds. This phenomenon is exemplified in incidents like flash crashes — i.e., sudden and drastic market fluctuations driven by the high-speed responses of automated trading systems.
Rights infringement. The pervasive application of AI in mass surveillance and extensive data collection practices is increasingly encroaching upon basic human rights. Privacy breaches have become alarmingly routine as AI systems gather and analyze personal data on an unprecedented scale. This constant monitoring undermines the right to privacy, a cornerstone of individual freedom. Furthermore, these surveillance mechanisms exert a chilling effect on freedom of expression. Individuals, aware of the omnipresent AI-driven surveillance, may self-censor or refrain from expressing dissenting opinions, leading to a stifling of public discourse and democratic engagement. The situation is compounded by AI systems that enable discriminatory profiling and unwarranted scrutiny of individuals. Such practices, often lacking transparency and accountability, lead to systemic unjust treatment and exacerbate existing societal inequalities at scale.
In the perfect storm MISTER Scenario, a series of interconnected AI-induced risks coalesce into a catastrophic sequence of events, each exacerbating the next, leading to a potential existential crisis for humanity.
The AI x-catastrophe unfolds with a devastating AI-driven cyberattack simultaneously targeting critical power grids across three continents. This orchestrated attack is the tipping point, a culmination of the escalating cybersecurity threats. The resultant continent-wide blackouts cause immediate and widespread chaos, disrupting essential services and plunging billions into darkness. The blackouts trigger a domino effect, causing major economic crashes. Financial markets, already destabilized by AI-induced manipulations, collapse under the strain. The economic fallout rapidly fuels societal unrest, with widespread protests and riots in response to the failing systems.
Amidst this chaos and darkness, the seeds of distrust sown by AI-manipulated media, deepfakes, and targeted disinformation campaigns, which had been proliferating prior to the blackouts, begin to bear fruit. These divisive narratives, deeply entrenched in public consciousness, exacerbate social divides and impede efforts to restore stability and order. The blackout acts as a catalyst, propelling these latent tensions into active, widespread civil unrest. Simultaneously, the crisis exposes and amplifies previously minor inefficiencies and errors in AI systems, which become more pronounced due to the volatile market dynamics, regulatory upheaval, and ongoing algorithmic adjustments. These AI system failures extend their impact across various critical infrastructures, including healthcare and communication networks, further amplifying the societal disruption.
The causal impact of AI inefficiencies limited to each subsystem, each seemingly non-existential in isolation, begins to accumulate dynamically and gives rise to compounded systemic impact, leading to disastrous global impacts. The convergence of these catastrophic events — multiple cyberattacks, manipulation, systemic eroded trust, economic destabilization, and rights infringements — leads to a state of global dysfunction. The capacity for a coordinated global response becomes critically undermined, as nations grapple with internal crises and widespread infrastructural breakdowns. Regional conflicts escalate into larger wars. Nations or non-state actors, driven by desperation or opportunism, engage in aggressive military actions, potentially leveraging AI technologies in warfare without legal constraints.
In this scenario, the x-catastrophe emerges from the synergistic failure of systems critical to the functioning and survival of human civilization. The simultaneous and compounded nature of these crises creates a perfect storm situation where not only is recovery extremely challenging, but the potential for irreversible collapse is a stark reality.
Before discussing the ramifications of the accumulative AI x-risk hypothesis, let’s first consider and address two potential objections.
7 Objections and replies
Objection 1: The societal breakdown or accumulative civilizational collapse, as it is characterized, does not equate to the x-catastrophes often envisioned in AI risk discussions, such as an ASI power-seeking entity deliberately aims to destruct humanity. Historically, civilizational collapses have happened and, while significant, have not equated to human extinction or unrecoverable collapse.
Reply 1: This objection, while historically grounded, may not fully consider the unique AI x-catastrophic threats posed in our modern, interconnected global civilization. The modern world functions as a tightly interconnected global system, far more integrated than any past civilization. The COVID-19 pandemic serves as a contemporary example, where a health crisis originating in one region quickly escalated into a global emergency, disrupting economies, supply chains, and everyday life worldwide. The perfect storm MISTER example illustrates how a collapse in one sector or region can have rapid, global impacts in today’s interconnected world, a phenomenon not seen in past civilizational collapses.
Objection 2: The accumulative AI x-risk model is too complex and unpredictable. Tracking and predicting the cumulative effects of various smaller AI-induced disruptions over time may be impractical or impossible, thus rendering this model less useful for x-risk assessment and mitigation.
Reply 2: The complexity inherent in the accumulative AI x-risk model is actually a key advantage. It promotes a comprehensive understanding of the diverse impacts of AI across societal, economic, and ecological domains. By acknowledging the interdependencies and potential for cascading effects within these systems, we can increase resilience and adaptability against x-catastrophic disruptions, leading to the development of more robust long-term strategies.
While the paperclip maximizer example offers a simple and elegant model, it deeply overlooks the nuanced realities of AI’s interactions with complex global systems. The accumulative model, in contrast, acknowledges this complexity and provides a more realistic portrayal of potential risks. It emphasizes detailed, empirically-grounded analysis, which is crucial for identifying critical areas sensitive to existential risks.
Future efforts should aim to further validate and refine the accumulative hypothesis through system dynamic simulations (e.g., Karnopp et al. (2012)). These simulations would offer tangible insights into the complex interplay of AI-induced risks, enhancing our understanding and ability to model these interactions accurately. Although this paper does not resolve all the questions surrounding the accumulative model, it shows the need for ongoing research and deeper exploration in the effective management of AI x-risks. The approach towards a more nuanced understanding of AI’s potential impact on our world is both necessary and urgent.
8 Implications
In this section, I briefly discuss five governance implications of this expanded conception of AI x-risks.
8.1 Consolidated approach to AI risk governance
Recognizing the distinct nature of decisive and accumulative AI x-risks calls for a holistic evaluation of how we govern AI risks.
The governance of decisive ASI x-risks may be best served by centralized and stringent control measures. These controls aim to quickly address and mitigate the immediate and widespread consequences that such particular risks could unleash. For instance, stringent regulation of ASI research and development is needed. This might include international agreements akin to the Nuclear Non-Proliferation Treaty, ensuring that advancements towards ASI are closely monitored and controlled. Rapid-response protocols, similar to those used in handling biochemical threats, could be implemented.
In the event of an ASI-related crisis, a centralized authority could quickly mobilize resources and expertise to mitigate the threat. On the other hand, accumulative AI x-risks, which develop gradually and are often less perceptible in the short term, calls for decentralized, adaptable governance. This could involve a network of local and regional AI oversight bodies that monitor the impacts of AI in their respective areas, akin to environmental protection agencies that oversee local environmental concerns.
The unification of different domains in AI risk governance not only streamlines efforts but also avoids the unnecessary costs associated with rebranding. AI development is a gradual and continuous process, and similarly, the development of tools for detecting and mitigating AI risks should also be progressive and sustained. The integrated approach prevents the redundant effort of ’reinventing the wheel’ for each distinct aspect of AI x-risk mitigation.
To clarify, the consolidated approach to AI risk governance does not necessitate a direct link between every project and each facet of social risks to x-risks. Rather, it advocates for a meta-view or a macro-level perspective. This perspective facilitates a comprehensive understanding of how various projects, while not necessarily directly interlinked, collectively contribute to the overarching objective of mitigating AI x-risks. In light of this discussion, I propose to establish a dynamic ”risk contribution checklist.” This checklist, designed to be both flexible and regularly updated, would serve as a tool for researchers and institutions engaged in AI development. It would enable a coordinated approach to risk mitigation, ensuring that diverse projects align with and contribute effectively to the broader goal of managing AI existential risks.
8.2 Harmonization of ethical and existential risks
Traditionally, particularly in the Anglo-American context, there has been a tendency to treat communities focusing on ethical risks and those concerned with x-risks as distinct (Section 2). However, the accumulative AI x-risk hypotheses challenge this separation, indicating that such a dichotomy is epistemically unsound in the context of AI risks.
In addition, the distinction between decisive and accumulative AI x-risks suggests a potential realignment of AI x-safety research priorities. While investigating catastrophic failure modes associated with ASI remains crucial, there is a growing need to allocate comparable attention and resources to the exploration of more subtle, long-term existential effects of AI via ethical risks. These effects, often cumulative in nature, require thorough study to understand their full impact and implications. Neglecting the accumulative aspect of AI x-risks would be a critical oversight. Failing to recognize and investigate these gradual and accumulative impacts could leave us unprepared for the nuanced existential challenges.
Moreover, the harmonization facilitates the transfer of risk mitigation strategies across different projects. A prime example of this is the evolution of interpretable models to mitigate social AI risks into mechanistic interpretability to mitigate AI x-risks. This progression demonstrates how methodologies initially developed in one domain can be adapted and enhanced for application in another, thereby ensuring a more cohesive and efficient approach to managing AI risks.
8.3 Short-long-term risk management
In the literature, there has been an implicit assumption that it is plausible to separate short-term (ethical) and long-term (existential) AI risks. As I have suggested, upon closer examination, these lines blur, revealing a complex interplay of concerns that transcend traditional time frames. The immediate risks of advanced AI systems is not solely confined to the present but also encompasses the rapidly approaching future, indicating the interconnected nature of these risks. As such, the structure of AI risk-related problems indeed represents a complex blend of immediate, intermediate, and long-term considerations, challenging our conventional understanding of risk and time.
Recognising the accumulative emergence of x-risks as a result of mismanagement, accumulation, or contingency path through ethical risks allows us to plan for schedules that address simultaneously both short-term and long-term risk mitigations and emphasise the importance of ongoing monitoring and evaluation. It is interesting to explore how current AI risk management frameworks (Baryannis et al., 2019; Tabassi, 2023) adapt to the accumulative and decisive hypothesis in this paper.
8.4 Enhancing Understanding and Awareness of AI Risks
The critical link between non-existential social risks and x-risks shows the importance of elevating public awareness and education regarding the radical impacts of AI. This effort is geared towards empowering individuals and communities to participate meaningfully in conversations surrounding responsible AI development.
Decisive x-risks, characterized by their immediate and striking effects, has tended to capture public attention. They resonate with cognitive biases, tap into inherent fears, or are amplified by sensationalist media portrayals. This dynamic may result in a distorted perception of AI x-risks, where the immediacy and drama of these risks eclipse the more subtle and gradually developing accumulative x-risks. In contrast, the accumulative risks, owing to their incremental nature and less obvious long-term impacts, often do not generate the same level of public interest or urgency. They typically do not provoke the instant emotional reaction seen with decisive x-risks, potentially leading to delayed public recognition and response until these risks escalate to critical levels.
Therefore, it is important to steer the conversation on AI risks away from purely dystopian or utopian narratives. While such narratives may be engaging, they frequently lack the necessary depth and complexity for constructive discussion and effective risk management. A more nuanced and informed approach to discussing AI risks is crucial for fostering a balanced understanding and preparing for the future with AI.
8.5 Mitigating x-catastrophes by limited connectivity
Finally, the distinction between two types of AI x-risks explores the strategy of mitigating decisive AI-related x-catastrophes by deliberately avoiding universal connectivity to ASI. Understanding the risk landscapes through systems analysis reveals the potential benefits of establishing multiple, critical nodes. This strategic decentralization serves as a protective measure, enhancing global resilience against the spread of catastrophic events, particularly those instigated by ASI.
If the paper clip maximizer is not expansively connected to humanity, the potential for decisive AI x-risk substantially diminishes. Consequently, the discussion so far advocates for a cautious stance against the globalization of ASI.
9 Concluding remarks
I began by asking whether the conception of x-risk from AI systems relate to the conception of non-existential social risks. Conventional discussions of AI x-risks often respond negatively to this question, and overlook an alternative way by which AI could bring about x-catastrophes. Against this background, I formulated an alternative view of AI x-risk. I used systems analysis perspective to differentiate between the two conceptions of AI x-risks in relation to two distinct causal paths that could have AI x-catastrophes as their endpoint.
The two type of causal paths differ based on the way they instantiate three path features: non-linearity, connectedness, and dynamic system adaptation. Causal pathways that feature humanity cascade effect of ASI, super-connectedness of infrastructures to ASI, and system adaptation to unidirectional feedback loops, can end in decisive ASI x-catastrophes. Causal pathways that feature local causality with compounded systemic impact, selective-connectedness of infrastructures, and dynamic system adaptation to multidirectional feedback loops, can end in accumulative AI x-catastrophes. So x-risks should be about those events that occur in pathways and mitigation strategies targetted at them. I used a scenario termed perfect storm MISTER to propose how the accumulative model can give rise to AI x-catastrophes. I also discussed five governance implications of my proposed expanded conception of AI x-risks.
Going forward, I think there is no inherent reason to consider that the accumulative hypothesis is any less likely than the decisive view. The need to further substantiate the accumulative hypothesis is apparent. This could be effectively addressed through the development of computational simulations using a system dynamics approach, aligning with the perspective I have established in analyzing the two types of AI existential risks. Finally, my discussion has so far only differentiated between two types of AI x-risks. The possibility of additional categories remains an open question. The implications of my arguments for assessing the probabilities of catastrophic outcomes and existential risks are significant and warrant further exploration in future research.

### v3

@@
 While the former envisions an overt AI takeover pathway, characterized by scenarios like uncontrollable superintelligence, the latter suggests a different causal pathway to existential catastrophes.
-This involves a gradual accumulation of critical AI-induced threats such as severe vulnerabilities and systemic erosion of econopolitical structures.
-The accumulative hypothesis suggests a boiling frog scenario where incremental AI risks slowly converge, undermining resilience until a triggering event results in irreversible collapse.
+This involves a gradual accumulation of critical AI-induced threats such as severe vulnerabilities and systemic erosion of economic and political structures.
+The accumulative hypothesis suggests a boiling frog scenario where incremental AI risks slowly converge, undermining societal resilience until a triggering event results in irreversible collapse.
 Through systems analysis, this paper examines the distinct assumptions differentiating these two hypotheses.
-It is then argued that the accumulative view reconciles seemingly incompatible perspectives on AI risks.
-The implications of differentiating between these causal pathways — the decisive and the accumulative — for the governance of AI risks as well as long-term AI safety are discussed.
+It is then argued that the accumulative view can reconcile seemingly incompatible perspectives on AI risks.
+The implications of differentiating between these causal pathways — the decisive and the accumulative — for the governance of AI as well as long-term AI safety are discussed.
+Forthcoming in Philosophical Studies
 1 Introduction
 Recent advances in machine learning have sparked intense debate about the existential risks (x-risks) associated with Artificial intelligence (AI) systems.
+[^1] For a review of recent narratives about AI and future, see Gilardi et al.
+(2024).
 Central to this debate is a concern about the potential pathways through which AI could cause existential catastrophes.
-In direct response to this concern, this paper aims to explore: How should we effectively conceptualize AI x-risks in the context of various types of causal pathways leading to AI-induced existential catastrophes?
-Conventional discourse on AI existential catastrophes typically portrays them as sudden, decisive events, often triggered by artificial general intelligence or artificial superintelligence (e.g., Bostrom (2013) and Ord (2020a)).
-Contrasting this traditional decisive viewpoint, this paper introduces the accumulative AI x-risk hypothesis as an alternative lens.
-The accumulative hypothesis posits that AI x-risks do not exclusively materialize as high-magnitude events initiated by artificial general intelligence or artificial superintelligence.
-Instead, they can accumulate incrementally through a series of smaller yet consequential AI-induced disruptions over prolonged periods.
-These disruptions interact synergistically, gradually undermining systemic resilience, ultimately leading to a critical point where a stressor event could result in an irreversible collapse.
-This paper defends and develops this accumulative perspective which is not adequately represented or robustly defended in the philosophical literature.
-I argue that a thorough and nuanced understanding of AI x-risks necessitates the integration of this accumulative hypothesis into both qualitative and quantitative x-risk studies.
-Recognizing the accumulative nature of AI x-risks is critical for the governance of AI risks and the development of long-term AI safety initiatives.
-This paper hence suggests a reevaluation of existing approaches to AI x-risks in light of the accumulative hypothesis and encourages a more diverse perspective to AI x-risk management.
+In direct response to this concern, this paper explores: What are the distinct types of causal pathways through which AI systems could trigger existential catastrophes?
+Conventional discourse on AI existential catastrophes typically portrays them as sudden, decisive events, often triggered by artificial general or super intelligence (Bostrom, 2013) or extremely powerful (non-general) AI (Carlsmith, 2022).
+Contrasting this conventional decisive viewpoint, this paper introduces accumulative AI x-risk hypothesis as an alternative lens.
+The accumulative hypothesis posits that AI x-risks do not exclusively materialize as decisive, high-magnitude global events initiated by extremely powerful AI such as artificial general or super intelligence.
+Instead, locally significant AI-driven disruptions can accumulate and interact over time, progressively weakening the resilience of critical societal systems, from democratic institutions and economic markets to social trust networks.
+When these systems become sufficiently fragile, a modest perturbation could trigger cascading failures that propagate through the interdependence of these systems.
+The failures amplify and reinforce each other through network effects and feedback loops, potentially leading to an irreversible civilizational collapse.
+This paper develops an accumulative perspective on AI existential risk, by examining how multiple AI risks could compound and cascade over time to bring about an AI-generated existential catastrophe.
+Through applying systems analysis — an approach not typically used in AI x-risk scenario development — I defend the significance of accumulative AI x-risk, and consequently argue for a fundamental reconceptualization of AI x-risk governance, if such risks are to be governed effectively.
 2 AI x-risk: preliminaries
-Existential risks are risks associated with existential catastrophic events.
-These events result in either the extinction of humanity or an unrecoverable decline in humanity’s potential to thrive (Ord, 2020a).
+2.1 Concepts of risk
+At most basic, risk relates to some characterization of uncertainty about potential (adverse) outcomes (Dean, 1998; Hansson, 2010; Aven, 2012).
+According to the ISO 31000 standard (International Organization for Standardization, 2018), risk is defined as “the effect of uncertainty on objectives.” The Society for Risk Analysis (Aven et al., 2018) defines it as “uncertainty about and severity of the consequences of an activity,” while the U.S.
+Environmental Protection Agency (U.S.
+Environmental Protection Agency, 2024) defines risk as predicting “the probability, nature, and magnitude of the adverse effects that might occur.”
+More concretely, risk has been analyzed through at least four distinct, though not mutually exclusive, interpretations.
+First, risk has been used to mean an unwanted event that may occur (Carlsmith, 2022).
+For instance, “AI systems exhibiting power-seeking behavior pose a major existential risk to humanity.” Second, risk has been used to denote the cause of an unwanted event that may occur (Weidinger et al., 2022).
+For example, “Insufficient testing of generative AI models is a significant risk in AI deployment.” Third, risk has been used to mean the probability of an unwanted event that may occur (Lowrance, 1976).
+An example is “The risk that a large language model will generate incorrect information in response to user queries is approximately 5%.” Fourth, risk has been used to mean a statistical expectation value — the product of probability and consequence — of an unwanted event that may occur (United Nations International Strategy for Disaster Reduction, 2009).
+For instance, “The risk of LLM hallucination in a medical context can be calculated by multiplying the probability of giving incorrect medical advice (1%) by the average cost of medical liability claims (\$500,000), yielding an expected cost of \$5,000 per consultation.”
+The particular focus of this paper is existential risks from AI.
+Existential risks (x-risks) refer to the potential for outcomes that would result in the extinction of humanity or an unrecoverable decline in humanity’s potential to thrive (Ord, 2020a).
 [^2] Some definitions, such as the one proposed by Bostrom (2013, p.
 15), broaden the scope of existential threats to include not only human life but all sentient beings: “An existential risk is one that threatens the premature extinction of Earth-originating intelligent life or the permanent and drastic destruction of its potential for desirable future development.” Despite this broader definition, this paper deliberately employs the term ”humanity” as the focal subject of AI-induced existential catastrophes (Ord, 2020a).
-The choice is pragmatic and is motivated by the predominant narrative in AI x-risk scholarly literature and public discourse, which has classically concentrated on the catastrophic impacts on humanity and human civilization.
-Additionally, this choice aligns with leading AI research laboratories such as Google DeepMind or OpenAI regarding the development of AGI or ASI for all of humanity.
-This terminological choice hence does not undermine the significance of non-human sentient beings.
+This choice is pragmatic and reflects the predominant narrative in AI x-risk scholarly literature and public discourse, which has traditionally concentrated on catastrophic impacts on humanity and human civilization.
+Additionally, this choice aligns with the stated missions of leading AI research companies like Google DeepMind and OpenAI in their pursuit of AGI or ASI development ”for the benefit of humanity.” My choice of terminology, however, does not diminish the moral significance of non-human sentient beings.
 Existential catastrophes are a class of potential events that may originate from natural causes, such as a supervolcanic eruption; anthropogenic sources, like nuclear conflict; or emerging threats, such as misaligned artificial superintelligence.
-[^2] For a historical analysis of existential risks, see Torres (2023).
-This paper concentrates on existential catastrophes induced by AI (AI x-catastrophes) and the associated AI x-risks.
+[^3] For a historical examination of existential risks, see Torres (2023b).
+This paper concentrates on existential catastrophes induced by AI (AI x-catastrophes) and their associated risks.
+In AI x-risk studies and discourse, researchers employ either of the qualitative and quantitative interpretations described above, each highlighting specific information about x-risks but also presenting specific limitations (Tonn and Stiefel, 2013; Beard et al., 2020).
+[^4] Quantitative methods for estimating AI x-risk require employing statistical models and subjective probability analysis.
+These methods are valuable for a seemingly structured way of estimating risks.
+However, they face limitations due to the reliance on available data, which can be scarce or unreliable for unprecedented risks, and the difficulty of robustly approximating the multifaceted nature of existential threats from AI in numerical terms.
+Qualitative methods involve non-numerical analyses such as scenario development, expert interviews, and ethical deliberations (Technology and Science Insights and Foresight, 2023).
+The qualitative methods are particularly useful in exploring the nuanced, complex, and often speculative aspects of AI risks, especially in areas where empirical data is scarce.
+This paper aims to maintain neutrality between these different interpretations of x-risk, drawing on a broad range of perspectives to analyze AI catastrophes and their associated risks.
+While I employ the causal interpretation of AI existential risks for illustrative purposes, alternative interpretations could remain equally valid.
+The probabilistic and statistical interpretations of x-risk warrant their own detailed investigation elsewhere.
 The conventional discourse on AI x-catastrophes portrays them as decisive, large-scale events caused by highly advanced AI systems, often referred to as artificial general intelligence (AGI) or artificial superintelligence (ASI).
-[^3] In AI lexicon, AGI roughly represents a capability level on par with human intelligence, while ASI denotes capabilities surpassing human intelligence in achieving objectives.
-Typically, it is argued that AI x-risks are caused by AGI or ASI (i.e., Bostrom (2014); Ord (2020a, b)).
-Another perspective suggests that AI x-catastrophes might arise from AI systems lacking complete general or super intelligence, yet possessing specific extremely powerful capabilities (e.g., Carlsmith (2022)).
-The preference to tie x-risks to AGI or ASI, rather than extremely capable but narrow AI systems, in this paper is motivated by the following consideration: the x-risks from AGI or ASI have been employed by influential voices in the academic field of x-risk studies as well as pioneering AI research labs such as OpenAI.
-The decisive nature of x-risks according to the conventional discussions equally holds if we replace the term AGI or ASI with extremely dangerous capable yet narrow AI that are not generally or superly intelligent, but have very narrow-focused uncontrollable capabilities.
-The terminological choice for AGI or ASI x-risks, therefore, does not affect the characterization of the conventional view on AI x-risk in this paper.
-The idea that advanced machines can pose significant risks is not fully novel and has historical antecedents.
+[^5] In AI lexicon, AGI represents capabilities comparable to human intelligence, while ASI denotes capabilities surpassing human intelligence.
+Traditional arguments typically link x-risks primarily to AGI or ASI (e.g., Bostrom (2014); Ord (2020a, b)), though some suggest that AI x-risks could also emerge from AI systems with extreme but narrow capabilities (Carlsmith (2022)).
+This paper’s focus on AGI/ASI aligns with influential voices in x-risk studies such as major AGI research companies.
+However, my core argument about the nature of x-risks holds whether we consider AGI/ASI or narrower AI systems with extreme, uncontrollable capabilities in specific domains.
+Thus, this terminological choice does not affect my characterization of conventional views on AI x-risk.
+The idea that advanced machines can pose significant risks is not novel and has historical antecedents.
+[^6] For a comprehensive historical review, see Torres (2023a).
 Samuel Butler (1863, p.185), a novelist and literary critic, alluded to the possibility of machines dominating humanity.
@@
 Similarly, the mathematician Irving J.
-Good (1966) expressed concerns about the creation of ultraintelligent machines.
+Good (1966) expressed concerns about the creation of “ultraintelligent machines.”
 More recently, philosophers like Nick Bostrom (2002, p.7) drew systematic attention to the existential threats posed by ASI: “When we create the first superintelligent entity, we might make a mistake and give it goals that lead it to annihilate humankind, assuming its enormous intellectual advantage gives it the power to do so.” [^7] While Bostrom’s work has systematized the discussion of x-risks from ASI, the original post-2000 discussions trace back to Eliezer Yudkowsky’s views concerning AGI, ASI, and their associated x-risks.
 Ben Goertzel (2015) explores the early conceptual evolution of these topics, tracing their roots to Yudkowsky’s initial informal explorations.
-I primarily reference Bostrom as a representative figure who brought systematic and philosophical depth to concepts that Yudkowsky and others initially introduced and examined in speculative media and blog posts.
+In this paper, I primarily reference Nick Bostrom as a representative figure who brought systematic and philosophical depth to concepts that Yudkowsky and others initially introduced and examined in speculative media and blog posts.
 Computer scientists such as Stuart Russell (2019) and physicists like Max Tegmark (2018) echoed similar concerns, stressing the x-risks of ASI beyond human control.
 Such views about ASI x-risks frequently hinge on two key theses: orthogonality and instrumental convergence (Bostrom, 2012, 2014).
-The orthogonality thesis posits an AI system’s intelligence and its goals are independent variables.
-This implies that an ASI could be designed with any number of objectives, benevolent or otherwise, irrespective of its intellectual capacity.
-That is, an AI system can be superintelligent but possess x-catastrophic goals as readily as it could have beneficial ones.
-The instrumental convergence thesis proposes that ASI would seek self-preservation and goal achievement by any means necessary.
-This could include removing obstacles like humans.
-Two hypothetical scenarios illustrate conventional models of causal pathway derived from the dual theses, which potentially lead to x-catastrophic outcomes from ASI.
-In a thought experiment by Bostrom (2003), an ASI is given the goal of maximizing paper clip production.
-This AI could quickly realize that human existence is counterproductive to its goal.
-Humans pose a threat as they might deactivate the AI, resulting in fewer paper clips.
-Additionally, the AI might recognize that human bodies contain atoms that could be repurposed into paper clips.
-The ASI’s optimal future then involves a world abundant with paper clips but devoid of humans.
-In a second thought experiment, Russell and Norvig (2010, p.1039) reference a hypothetical scenario attributed to Marvin Minsky, a pioneer in AI research, where an advanced AI, tasked with solving the Riemann hypothesis, could decide to commandeer Earth’s resources to construct supercomputers, thereby aiding its objective and endangering humanity’s existence.
-The two hypothetical examples illustrate how ASI could independently pursue its goals to the detriment of humanity, and could deliberately or accidentally cause x-catastrophes, particularly when humans are unable to effectively steer the superintelligence’s behavior.
-Both experiments have in common the following causal pathway: a superintelligent reward-maximizer paired with a goal system that gives rewards for achieving existentially-catastrophic goals.
-On the conventional view sketched above, the path to the occurrence of AI x-catastrophes is via conclusive actions of ASI.
+[^8] For a critical discussion of the orthogonality and instrumental convergence theses, see Müller and Cannon (2022).
+The orthogonality thesis posits that an advanced AI system’s intelligence level and its final goals are orthogonal.
+[^9] See Bostrom (2012, p.73) and Bostrom (2014, p.107).
+This implies that an AI system could have any combination of final goals, beneficial or harmful, and intelligent capabilities.
+The instrumental convergence thesis holds that diverse final goals often have similar instrumental sub-goals — like self-preservation or resource acquisition — as these sub-goals are useful for achieving almost any final objective.
+An ASI might therefore pursue actions harmful to humanity not out of malice, but as instrumental steps toward achieving its programmed objectives.
+Two hypothetical scenarios illustrate how conventional models of ASI development could lead to x-catastrophic outcomes.
+In a thought experiment popularized by Bostrom (2003), an ASI is given the seemingly innocuous and simple goal of maximizing paperclip production.
+Even with this simple objective, the ASI could pursue instrumental sub-goals: it might eliminate humans to prevent deactivation (self-preservation) or convert their bodies into paperclips (resource acquisition).
+Consequently, the ASI’s optimal future could become one abundant with paperclips but devoid of humans — not because intelligence necessitates this goal, but because instrumental sub-goals are (putatively) rational steps toward satisfying its given optimization objective.
+That is, the paperclip maximizer illustrates how an AI with an apparently harmless goal could pose an x-risk through the rational pursuit of instrumental sub-goals like resource acquisition and self-preservation.
+[^10] The paperclip maximizer is characterized by some as a “Squiggle minimizer”: a highly intelligent optimizer pursuing goals alien to human values could inadvertently destroy humanity by consuming vital resources in pursuit of its objectives.
+See LessWrong (2024).
+In a structurally similar thought experiment, Russell and Norvig (2010, p.1039) describe a scenario attributed to Marvin Minsky: an advanced AI tasked with proving the Riemann hypothesis might appropriate Earth’s resources to build supercomputers, endangering humanity in pursuit of a mathematical proof.
+Like the paperclip maximizer, this example is supposed to illustrate how an advanced AI pursuing a benign goal could pose existential risks to humanity.
+Both scenarios demonstrate a key pattern: an advanced AI system optimizing for a specific goal could rationally pursue sub-goals which are catastrophic to humanity, not from malice but as instrumental steps toward its optimization goal.
+2.2 Decisive AI x-risk
+The conventional view, sketched above, frames AI x-catastrophes arising through decisive actions of ASI.
 Toby Ord (2020a, p.
-20), a notable figure in x-risk studies, explicitly endorses the decisive character of x-catastrophes: “I take on the usual sense of catastrophe as a single decisive event rather than any combination of events that is bad in sum.
-A true existential catastrophe must by its very nature be the decisive moment of human history the point where we failed.” Given that AI x-risks concern the potential for AI x-catastrophes, the conventional viewpoint suggests that AI x-risks are about sudden, cataclysmic events that could either eradicate humanity or significantly curtail its potential.
-The core assumption underlying this viewpoint can be articulated as the decisive ASI x-risk hypothesis:
+20), a prominent x-risk scholar, explicitly endorses the decisive character of x-catastrophes: “I take on the usual sense of catastrophe as a single decisive event rather than any combination of events that is bad in sum.
+A true existential catastrophe must by its very nature be the decisive moment of human history the point where we failed.” This conventional framing suggests that AI x-risks manifest as sudden, cataclysmic events that either eradicate humanity or irreversibly curtail its potential.
+I articulate this perspective in terms of the decisive ASI x-risk hypothesis:
 Decisive ASI x-risk hypothesis: x-risks from ASI concern the possibility of abrupt large-scale events that lead to humanity’s extinction or cause an unrecoverable decline in its potential.
-Bostrom’s portrayal of ASI incorporating extremely dangerous goals leading to the annihilation of humanity points at a conclusive event with extremely dire consequences.
-Similarly, Ord’s characterization of a true x-catastrophe as a single, defining event illustrates the notion of AI x-risk being about shockingly high-impact events.
-The decisive AI x-risk, in this context, is the expected uncertainty of the occurrence of such conclusive events as catalysts for x-catastrophic outcomes.
-[^5] Typically, the term ”risk” is quantitatively defined in terms of likelihood and impact.
-This quantitative approach involves assessing the probability of a specific event occurring and quantifying the potential consequences.
-However, in x-risk studies, researchers employ both qualitative and quantitative methods, each offering unique insights but also presenting specific limitations (Tonn and Stiefel, 2013; Beard et al., 2020).
-Quantitative methods for estimating AI x-risk require employing statistical models and subjective probability analysis.
-This approach is valuable for a seemingly structured way of estimating risks.
-However, it faces limitations due to the reliance on available data, which can be scarce or unreliable for unprecedented risks, and the difficulty of robustly encapsulating the multifaceted nature of existential threats in numerical terms.
-Qualitative methods involve non-numerical analyses such as scenario development, expert interviews, and ethical deliberations (Technology and Science Insights and Foresight, 2023).
-This approach is particularly useful in exploring the nuanced, complex, and often speculative aspects of AI risks, especially in areas where empirical data is scarce.
-However, it lacks numerical precision.
-This paper adopt a neutral stance regarding the quantitative or qualitative interpretation of AI x-risks.
-This neutrality allows me to draw on the broadest range of insights and perspectives in my analysis of AI x-catastrophes and their associated risks.
+Both Bostrom’s portrayal of ASI pursuing destructive goals and Ord’s characterization of existential catastrophes as singular, defining events emphasize the conventional view: AI x-risks manifest as sudden, decisive moments of overwhelming impact.
+The decisive AI x-risk, according to this framing, is the expected uncertainty of the occurrence of such conclusive events as catalysts for x-catastrophic outcomes.
 The decisive hypothesis, however, overlooks an alternative type of causal pathway leading to AI x-catastrophes.
@@
 In the rest of this paper, the terms ”AI ethical risk” and ”AI social risk” are used interchangeably.
+2.3 AI risks: existential versus social
 The prevailing discourse on categorizing AI risks distinguishes between AI x-risks and AI social risks as separate and distinct categories.
@@
 In a notable instance, Turing Award Laureate Geoffrey Hinton, who departed from his position at Google to openly discuss AI x-risks, emphasized during a recent interview that his concerns about existential risk ”are different” from Timnit Gebru’s concerns about AI ethical risks that are not ”existentially serious” (CNN, 2023).
-Typically, existential and social risks are demarcated along the lines of locality — i.e., the scope of risk — and severity — i.e., hurting one or all including future generations (Bostrom, 2013; Amodei et al., 2016).
-While this demarcation is insightful, there is a notable gap in exploring the relationship between x-risks and the evolution of ethical concerns in a substantial manner.
+Typically, existential and social risks are demarcated along the lines of locality — i.e., the scope of risk — and severity — i.e., whether impacts are recoverable and limited, or irreversible and catastrophic for (human) civilization (Bostrom, 2013; Amodei et al., 2016).
+While this demarcation is pragmatically insightful, there is a notable gap in exploring the relationship between x-risks and the evolution of ethical concerns in a substantial manner.
 [^13] This disconnect might stem from two key reasons.
 First, discussions about decisive x-risks historically have revolved around reinforcement learning and agent-environment models, concentrating on existential threats from agency-based models.
-Ethical risks, in contrast, embrace a more expansive conceptual framework.
+Ethical risks, in contrast, embrace a more expansive approach, covering non-reinforcement approaches to AI development.
 Second, prominent conventional voices in ASI x-risks, such as Bostrom, Ord, and Yudkowsky, often subscribe to worldviews like rationalism, effective altruism, or longtermism.
@@
 A thorough exploration of these distinctions, however, falls outside the scope of this paper.
-The social risks of AI systems have been extensively analyzed and studied, leading to the identification of six major risk categories: (1) lack of accountability and transparency about AIs and their owners, (2) representational and allocative unfairness and wrongful biases from AI predictions and automated decisions, (3) risk of the amplification of AI-generated misinformation, (4) threats to individual and group privacy, (5) malicious and dual use of AI, and (6) systemic risks impacting political, economic, and environmental stability.
-[^9] For lack of AI accountability and transparency, see Mitchell et al.
-(2019); Raji et al.
-(2020); Gebru et al.
-(2021); for representational and allocative unfairness and wrongful biases from AI predictions and automated decisions, see Buolamwini and Gebru (2018); Chouldechova and Roth (2020); Kleinberg et al.
-(2016); for the amplification risk of AI-generated misinformation, see Sharma et al.
-(2019); Quach (2020); Lin et al.
-(2021); for risks to individual and group privacy, see Dwork (2006); Tucker (2018); for risks of malicious and dual use, see Brundage et al.
-(2018); and for systemic risks impacting political economic, and environmental stability, see Korinek and Stiglitz (2018); Pashentsev (2021).
-For general taxonomies of the social risks of generative AI, see Weidinger et al.
-(2021); Bird et al.
-(2023); Barnett (2023).
-This paper posits that the gradual and cumulative progression of AI social risks delineates an alternative causal pathway to AI x-catastrophes:
-Accumulative AI x-risk hypothesis: AI x-risks accumulate through a series of lower-severity disruptions over extensive duration, collectively eroding systemic resilience until some stressor event triggers unrecoverable collapse.
-According to the alternative view, AI x-catastrophes could emerge not from a decisive event, but from the cumulative impact of multiple interconnected AI-induced adverse events over time.
-As compared to the decisive hypothesis, this accumulative hypothesis suggests a different causal pathway to AI x-catastrophe: a path wherein a succession of lower-severity, yet cumulatively significant, disruptions deeply erode the systemic resilience of the global system (which embodies humanity), radically disrupting socioeconomic and sociopolitical equilibrium.
-This weakened state potentially primes the system for an unrecoverable collapse, particularly when further stressed by external events.
-Figure 1 provides a schematic illustration that contrasts the two hypotheses: the decisive and accumulative models of AI x-risk escalation.
-The graph plots severity of AI risks against time, with two distinct trajectories.
-The decisive model, marked in blue, shows a sharp increase in risk severity over a short time span, indicating a sudden, high-impact event.
-In contrast, the accumulative model, represented in red, depicts a gradual, evolving increase in risk severity, reflecting a series of escalating but less severe events over time.
-A dashed line marks the x-catastrophe threshold, indicating the point at which either scenario leads to a potential x-catastrophe.
-0010101010202020203030303040404040505050500020202020404040406060606080808080100100100100X-catastrophe thresholdTimeSeverity of AI risksDecisive ASI x-riskAccumulative AI x-risk
+The social risks of AI systems have been analyzed across different domains — from language models and their multimodal variants (Bender et al., 2021; Weidinger et al., 2022; Bird et al., 2023) to recommender systems (Milano et al., 2020; Deldjoo et al., 2024), to name just a few examples.
+While a full exposition of these risks is beyond the scope of this paper and has been extensively discussed in the above references, here I provide a brief categorization.
+Manipulation and deception risks include AI systems that cause harm by manipulating human behavior through targeted or unwanted persuasion (Kasirzadeh and Evans, 2023; Carroll et al., 2024).
+Misinformation and disinformation risks arise from AI systems generating and amplifying false content at scale, enabling the spread of propaganda and undermining public trust and discourse (Sharma et al., 2019; Quach, 2020; Lin et al., 2021; Kay et al., 2024).
+Malicious use risks include the weaponization of AI systems for cyber attacks, the deployment of AI-enabled drones and other physical systems for attacks, as well as the automation of social engineering attacks (Brundage et al., 2018).
+Insecurity and information threat risks arise when AI systems reveal sensitive personal data or lead to an unintended disclosure of protected information (Carlini et al., 2021).
+Discrimination and hate speech risks manifest through biased AI systems perpetuating systemic inequalities or generating targeted harmful content (Buolamwini and Gebru, 2018; Obermeyer et al., 2019).
+Surveillance, rights infringement, and erosion of trust risks originate from AI-powered mass surveillance systems and persistent monitoring that could lead to loss of privacy (Dwork, 2006; Tucker, 2018) and loss of trust in ruling institutions (Nowotny, 2021).
+Environmental and socioeconomic risks result from massive energy consumption of AI training or displacement of workers through automation (Korinek and Stiglitz, 2018; Pashentsev, 2021).
+2.4 Accumulative AI x-risk
+As an alternative to the decisive AI x-risk hypothesis, the gradual and cumulative progression of critically significant social risks characterizes a different type of causal pathway to AI x-catastrophes:
+Accumulative AI x-risk hypothesis: AI x-risks result from the build-up of a series of smaller, lower-severity disruptions over time, collectively and gradually weakening systemic resilience until a triggering event causes unrecoverable collapse.
+According to this alternative view, AI x-catastrophes could emerge not from a decisive event, but from the cumulative impact of multiple interconnected AI-induced adverse events over time.
+As compared to the decisive hypothesis, the accumulative hypothesis suggests a different causal pathway to AI x-catastrophe: a path wherein a succession of lower-severity, yet cumulatively significant, disruptions deeply erode the systemic resilience of the global system, radically disrupting socioeconomic and sociopolitical equilibrium.
+This weakened state potentially primes the global system for an unrecoverable collapse, particularly when further stressed by external events.
+0010101010202020203030303040404040505050500020202020404040406060606080808080100100100100 ASI arrives Critical thresholdTimeSeverity of disruptionDecisive scenarioAccumulative scenario
 Figure 1: AI x-risk escalation: decisive and accumulative models
+Figure 1 provides a schematic illustration contrasting the two hypotheses: the decisive and accumulative models of AI x-risk escalation.
+The decisive scenario (blue line) represents the conventional view where a sudden catastrophic event — such as one caused by ASI — leads to rapid, irreversible consequences.
+In contrast, the accumulative scenario (red line) shows how existential catastrophes can arise through a gradual path, where multiple smaller disruptions interact and amplify over time through feedback loops and compounding effects.
+These accumulating disruptions gradually erode system stability, potentially crossing critical thresholds until a triggering event occurs.
+[^14] The time axis is purely illustrative and not tied to specific units, as the actual temporal development of these scenarios remains uncertain.
+The figure aims to capture qualitative differences in catastrophic events caused by AI rather than make specific predictions about timing or precise severity levels.
+The zigzag pattern in the accumulative scenario represents potential local recoveries, though the overall trend shows increasing systemic fragility over time.
 The gradual nature of the accumulative AI x-risk hypothesis can be likened to other global existential threats such as climate change or nuclear weapon proliferation.
-This is akin to the incremental rise in greenhouse gases contributing to climate change, where each individual emission seems minor, but collectively, they lead to significant and potentially irreversible changes in the Earth’s climate.
+[^15] Ord (2020b, p.
+28) acknowledges the decisive versus accumulative nature of global catastrophic events: “Nuclear weapons and climate change have striking similarities and contrasts.
+They both threaten humanity through major shifts in the Earth’s temperature, but in opposite directions.
+One burst in upon the scene as the product of an unpredictable scientific breakthrough; the other is the continuation of centuries-long scaling-up of old technologies.
+One poses a small risk of sudden and precipitous catastrophe; the other is a gradual, continuous process, with a delayed onset—where some level of catastrophe is assured and the major uncertainty lies in just how bad it will be.
+One involves a classifed military technology controlled by a handful of powerful actors; the other involves the aggregation of small effects from the choices of everyone in the world.” However, this point has not been made in the context of existential risks from AI systems.
+The accumulative perspective is structurally akin to the incremental rise in greenhouse gases contributing to climate change, where each individual emission seems minor, but collectively, they lead to significant and potentially irreversible changes in the Earth’s climate.
 Similarly, the steady build-up of nuclear weapons incrementally increases the risk of catastrophic conflict, with each addition subtly heightening global tensions.
-The key lesson here is that x-risks often stem not from sudden, dramatic events, but from the gradual accumulation of smaller incidents and decisions, which over time, can existentially escalate and diminish safety margins.
-In the rest of this paper, I investigate the nuanced assumptions differentiating the two hypotheses, employing a systems analysis perspective.
-More specifically, I differentiate between two types of causal pathways that could lead to AI x-catastrophes: the decisive and the accumulative pathways.
-Following this, the paper proposes an integrated approach to AI x-risk evaluation and governance.
-This approach bridges both the accumulative and decisive hypotheses, recognizing them as complementary causal pathways leading to AI x-catastrophes, and challenges the conventional view that AI x-risks and AI ethical risks are separate and unrelated.
-It posits that effective AI x-risk governance and long-term AI safety require a substantial focus on the development and accumulation of AI social risks, thus highlighting the necessity of addressing critical ethical risks in mitigating AI x-risks.
+The key lesson here is that x-risks could stem not from sudden, dramatic events, but from the gradual accumulation of smaller incidents or decisions, which over time, can existentially escalate and diminish safety margins.
+In the rest of this paper, I employ systems analysis to investigate two distinct causal pathways that could lead to AI x-catastrophes.
+The decisive pathway assumes direct, catastrophic failures triggered by a superintelligent system, while the accumulative pathway describes how systemic instabilities emerge from interactions between multiple AI-driven disruptions.
+This systems-analysis perspective allows us to formulate the key assumptions underlying each hypothesis and examine how different patterns of causation and connectivity could produce existential risks through fundamentally different mechanisms.
 Before proceeding further, a crucial clarification is in order.
 The emphasis of this paper on conventional discussions of AI x-risks does not imply that all discourse on AI x-risks has been confined to the decisive hypothesis.
-The discourse on AI x-risks is expanding, and an increasing number of scholars who do not necessarily endorse a decisive viewpoint are engaging with these issues (see, for example, Bucknall and Dori-Hacohen (2022), Hendrycks and Mazeika (2022), and Shevlane et al.
+This discourse is expanding, and an increasing number of scholars who do not necessarily endorse a decisive viewpoint are engaging with these issues (see, for example, Bucknall and Dori-Hacohen (2022), Hendrycks and Mazeika (2022), Shevlane et al.
+(2023), and Bales et al.
 (2024)).
@@
 This paper focuses on developing the accumulative perspective on AI x-risk which has not been adequately represented or robustly defended in the philosophical literature.
-It is my hope that this additional would facilitate a more unified and constructive dialogue about different kinds of AI risks.
-3 Two types of causal pathway
-Although it may appear obvious, a crucial yet frequently overlooked aspect in studies of AI x-risks and x-catastrophes is serious attention to their manifestation within a complex global system.
-In this system, human and AI components interact across a variety of domains, creating a multifaceted landscape for these risks and catastrophes to unfold.
-AI x-catastrophes and their associated AI x-risks hence represent the endpoints of particular causal pathways within this global interconnected system.
-Effectively identifying and mitigating AI x-risks, whether through qualitative or quantitative methodologies, demands an in-depth examination of the nature of these causal pathways.
-Building upon this observation, I employ a systems analysis perspective to shed light on two distinct categories of causal pathways leading to AI x-catastrophes.
-[^10] Rather than attempting to compile an exhaustive list of every potential route, the focus here is on highlighting and contrasting these two primary types to broaden our understanding of AI x-risks and guide a more inclusive and strategic approach to addressing their far-reaching consequences for humanity.
-But why is a systems analysis perspective advantageous in examining AI x-risk pathways, and what does it involve?
-This perspective, rooted in systems thinking as outlined by Meadows (2008), is instrumental in identifying the high-level characteristics of pathways in the global system that can lead to AI x-catastrophes.
-A system is an interconnected network of components that collectively exhibit specific behaviors or yield particular outcomes.
-Systems are characterized by their components’ interdependence and the boundaries that separate them from their surrounding environment.
-Systems thinking focuses on the interdependencies among various subsystems, uncovering a range of dynamic pathways.
-While it is not feasible to map all interrelations between these subsystems exhaustively, our focus primarily lies on meso-level subsystems such as political, socio-environmental, and economic infrastructures, recognizing their critical role in the broader dynamics of AI risks.
-Within the economic subsystem, AI’s development and deployment are deeply interwoven with market dynamics, investment trends, and corporate strategies.
-AI’s integration into economic infrastructure affects not only the allocation of resources for its research and development but also transforms the nature of economic activities and market operations.
+It is my hope that this addition would facilitate a more unified and constructive dialogue about different kinds of AI risk and their relations.
+3 Systems analysis for AI x-risk
+AI x-catastrophes and their associated risks appear within our complex global system.
+Understanding how these catastrophes could occur requires analyzing how various elements — humans, AIs, and organizations — interact across multiple domains within this system.
+Systems analysis, which provides conceptual and mathematical approaches for analyzing interactions between system components, enables us to trace how AI risks might propagate through these interconnected relationships.
+Systems analysis, pioneered by Von Bertalanffy (1968) and later advanced through influential works by Forrester (1971) and Meadows (2008), serves dual purpose for characterizing risks from AI.
+First, it provides an epistemic instrument for understanding how complex technological risks arise from interactions between multiple subsystems.
+Second, it offers a pragmatic instrument for identifying when and how to intervene in such systems to minimize the risk of undesired events.
+Systems analysis has proven valuable in understanding and taming various risks of global challenges — from the propagation of financial system shocks (Helbing, 2013) to the analysis of climate change tipping points (Steffen et al., 2018).
+But, what is a system?
+A system is a set of interconnected components whose interactions could produce emergent behaviors and outcomes.
+Three fundamental features define a system: its basic constitutive components, their interdependencies (how these elements influence each other), and its boundaries (what distinguishes this system from its environment).
+Systems analysis examines how phenomena evolve from initial perturbations that seed the process, through network propagation that spreads effects to compounding or cascading dynamics that amplify impacts, and finally to catastrophic transitions that transform entire systems.
+[^16] Catastrophe theory offers another relevant technical theory, as it explains how gradual changes in system parameters can trigger sudden, discontinuous transitions (Thom, 1974; Zeeman, 1977).
+While its mathematical formalism for analyzing tipping points and system transformations aligns well with our investigation of AI risks, its highly technical nature places a detailed application beyond this paper’s scope.
+Let me explain each in relation to AI risks.
+First, systems analysis provides a rigorous methodology for mapping how initial disturbances propagate through interconnected (AI) systems.
+For example, a software bug in one AI model might trigger failures in dependent systems, data corruption could cascade through shared training pipelines, or model misspecification might amplify errors across a network of automated decision systems.
+Second, the initial perturbations, however small, could establish or trigger transmission pathways through which effects can spread — much like how a small trading error can trigger a chain of automated responses in financial markets (Min and Borch, 2022), or how the failure of one power station can cascade through an electrical grid (Andersson et al., 2005).
+Third, systems analysis reveals cascade dynamics, where initial perturbations can trigger chains of events with amplifying effects.
+Buldyrev et al.
+(2010), for instance, show how local failures can overload other components, leading to consequential failures that spread through the network.
+In AI systems, these cascades can manifest in two key ways: through error cascades, where errors in one AI’s output become amplified as other systems build upon this flawed information, and through feedback loops, where initial disturbances cycle through the network, magnifying the original perturbation with each iteration.
+Fourth, systems analysis identifies potential catastrophic transitions at critical thresholds (Scheffer et al., 2009; Lenton et al., 2008).
+At these tipping points, seemingly small perturbations can trigger rapid, nonlinear changes in system-wide behavior.
+In AI systems, these might begin as minor fluctuations — such as subtle shifts in model behavior or isolated component failures — but once a threshold is crossed, they can fundamentally alter the entire network’s functioning and stability.
+Like a glass that shatters under increasing pressure, the system undergoes an abrupt transition from one state to another.
+The complex global system can be represented as a collection of interconnected subsystems.
+While a comprehensive mapping of all subsystems and their relationships is infeasible here, this analysis focuses on three critical meso-level subsystems — economic, political, and military.
+[^17] See World Economic Forum’s (2024) for an attempted mapping of the global network of interconnected subsystems and their associated risks.
+The economic subsystem, with its complex web of production, consumption, and exchange networks, could serve as a primary channel through which AI’s transformative effects propagate.
+As AI systems could increasingly substitute for routine cognitive tasks, they may reshape labor markets, particularly impacting middle-skilled service sectors.
+This transformation can also reverberate through capital markets via automated trading systems and risk assessment mechanisms, while fundamentally altering productivity dynamics through both labor augmentation and process automation.
+These changes can create feedback loops, amplifying economic disparities, as benefits accrue disproportionately to those with resources to adapt.
 [^18] See Korinek and Stiglitz (2018), Frey (2019), Agrawal et al.
-(2019), and Eloundou et al.
-(2023), among others.
-In the political subsystem, the direction of AI research and deployment is substantially shaped by national and international funding but also by establishing regulatory frameworks that dictate the limits of AI development and deployment.
-Additionally, the dynamics between various governments, international organizations, and global treaties contribute to a complicated geopolitical environment, steering AI research priorities.
-On the other hand, AI has profound implications for democratic elections and mass surveillance programs.
-AI technologies are increasingly used to manipulate public opinion, potentially undermining the integrity of elections through targeted misinformation campaigns, social media influence, and personalized political advertising, all driven by AI’s ability to analyze and predict human behavior and expression of desires.
-Advanced AI systems enable more sophisticated and pervasive surveillance, raising concerns about privacy, civil liberties, and the potential abuse of power.
-[^12] See Manheim and Kaplan (2019), Crawford (2021), and Madan and Ashok (2023), among others.
-In the military subsystem, AI’s incorporation into defense and surveillance subsystems highlights its strategic significance, drawing substantial research and funding attention.
-This military integration reflects AI’s escalating role in national security and defense strategies.
+(2019), Eloundou et al.
+(2023), and Acemoglu (2024) for varying perspectives on AI’s impacts across labor markets, productivity, and economic inequality.
+The political subsystem is typically linked to economic transformations.
+National and international funding decisions, regulatory frameworks, and geopolitical dynamics create a complex network of influences that direct AI research and deployment.
+Simultaneously, AI technologies are changing political processes themselves, enabling gradual manipulation of public opinion through targeted misinformation campaigns or personalized political advertising.
+[^19] See, for example, Collier et al.
+(2022) and Andrić and Kasirzadeh (2023).
+AI-infused surveillance and prediction technologies could create a self-reinforcing cycle: as governments increase surveillance, citizens resist these measures, leading to more intensive surveillance and control, ultimately weakening democratic institutions.
+[^20] See Manheim and Kaplan (2019), Crawford (2021), Madan and Ashok (2023), and Schaake (2024), among others.
+The military subsystem, deeply intertwined with both economic and political subsystems, represents another critical node.
+AI’s integration into defense and intelligence operations is core to national security strategies, while simultaneously influencing international relations.
+Military applications of AI could create feedback loops that affect both technological development priorities and geopolitical power dynamics.
 [^21] See Scharre (2018), Morgan et al.
 (2020), and Zuboff (2019), among others.
+The three subsystems do not operate in isolation but form a densely interconnected network where changes in one area compound through others, potentially in unexpected ways.
 Collectively, these subsystems, along with others, create the context within which AI x-catastrophes and their associated AI x-risks must be analyzed.
-To effectively discern the causal pathways that could lead to AI x-catastrophes and to scrutinize the risks associated with these paths, we need to understand key features of these pathways.
-The systems analysis perspective emphasizes three fundamental characteristics of pathway dynamics within a system: non-linearity, connectedness, and feedback loops.
-First, non-linear paths are intrinsic to complex systems, where small changes in one component can lead to disproportionate system-wide effects.
-This is often due to complex interactions among the system’s components.
-Second, connectedness is a critical element in systems analysis, highlighting how disturbances in one segment can cascade, leading to substantial changes in other areas.
-Third, feedback loops introduce cyclical dynamics where the outputs of processes consistently feed back into and influence subsequent inputs, shaping the behavior of the system.
-The following two sections delve into contrasting the characteristics of causal pathways leading to AI x-catastrophes as postulated by the decisive and accumulative hypotheses.
-This comparison is designed to show the distinct ways in which each hypothesis manifests the three aforementioned key features.
-4 Pathways to decisive ASI x-risk
-The decisive AI x-risk hypothesis considers scenarios where ASI could abruptly cause x-catastrophes.
-This hypothesis is anchored in specific manifestations of non-linearity, connectedness, and feedback loops, as follows: a rapid cascade effect within humanity triggered by ASI, the super-connectedness of various infrastructures with ASI, and the system’s adaptation to predominantly unidirectional feedback loops.
-($I_{D}$) A rapid cascade effect within humanity triggered by ASI.
-This characteristic aims to capture the extensive and profound impact of ASI, as it proliferates through or amplifies across various subsystems encompassing humanity.
-The connection to non-linearity lies in the manner in which ASI’s actions or presence could cause far-reaching and expansive consequences throughout the distribution of human civilization.
-For instance, in the paperclip maximizer scenario, the ASI’s ostensibly simple objective of maximizing paperclip production escalates into a global existential threat.
-This represents ASI’s capacity to redirect and dominate global resources by any means necessary, fulfilling its objective across the distribution of human civilization.
-($II_{D}$) Super-connectedness of infrastructures to ASI.
-This feature captures the idea that no significant subsystem can claim immunity from the influence of ASI’ actions on its path towards an AI x-catastrophe.
-This super-connectedness is the possible pervasive reach of ASI to all critical infrastructures.
-In scenarios like the paperclip maximizer, the ASI might use its access to subsystems (such as global manufacturing, global supply chains, and global energy grids) to further its paper clip production goal.
-This connectedness implies that the ASI can leverage all such necessary subsystems for the occurrence of an x-catastrophe.
-($III_{D}$) System adaptation to unidirectional feedback loops.
-The third feature of the decisive pathway is the dynamic adaptation of the global system, given unidirectional feedback loops that are both initiated and controlled by the ASI.
-These loops are mechanisms through which the ASI’s actions and decisions continuously reinforce and escalate its own objectives.
-The result is a system increasingly skewed towards the ASI’s goals, bypassing or overriding human-centric controls and checks.
-The absence or powerlessness of balancing feedback mechanisms leads to a situation where each loop intensifies the previous one, rapidly pushing the system towards extreme states aligned with ASI’s objectives.
-This dynamic evolution of the decisive causal pathway poses x-risk, as the ASI-driven loops can lead to situations that outstrip attempts at moderation, leading to what is often termed a singularity —a point of no return where the AI’s evolution is explosive and uncontrollable.
-In the paper clip maximizer scenario, the ASI self-optimizes for its task of producing paperclips.
-It becomes engaged in a self-perpetuating cycle of enhancement and resource acquisition, effectively becoming impervious to human intervention.
-5 Pathways to accumulative AI x-risk
-Recall the accumulative AI x-risk hypothesis, which serves as a counterpoint to the decisive perspective.
-The accumulative hypothesis suggests that AI x-risks may arise from an accumulation of lower-severity disruptions that interact significantly over an extended period, cumulatively weakening systemic resilience until a stressor event causes an unrecoverable collapse.
-This hypothesis is founded on distinct instantiations of non-linearity, connectedness, and feedback loops: local causality leading to compounded systemic impact, selective-connectedness among infrastructures, and the system’s adaptation to multidirectional feedback loops.
-Following the formulation of these pathway features, I will highlight more concrete manifestations of them in the ”perfect storm MISTER scenario” towards the end of this section.
-($I_{A}$) Local causality with compounded systemic impact.
-This feature highlights how clusters of localized AI impacts, even if minor at the outset, can aggregate and intensify across various subsystems.
-When these impacts reach a critical threshold, they pose a significant challenge to the overall systemic resilience.
-Unlike the immediate, broad-scale disruptions characteristic of the rapid cascade effect within humanity triggered by ASI, these local causal impacts evolve into significant threats through compounded changes within the global system via mechanisms such as ripple effect.
-The cumulative effect of substantial localized causal impacts, amplified by their frequency and extent of influence, can result in AI x-catastrophes.
-This gradual escalation entails a latency of broader-scope impact which implies a delay in the manifestation of effects where AI-induced alterations in specific areas radically transform broader societal behaviors and patterns.
-($II_{A}$) Selective-connectedness of infrastructures.
-The second pathway assumption is selective-connectedness of infrastructures to critical AIs.
-This stands in contrast to the broad, super-connectedness with ASI as outlined in assumption ($II_{D}$).
-Here, the focus is on a more discerning form of connectivity where not all substantial subsystems are uniformly linked to superintelligent AI.
-This selective-connectedness posits that certain critical AI-driven subsystems within the global system, while not immediately linked to widespread catastrophic outcomes, can, over time, induce significant systemic transformations.
-These changes can escalate to x-catastrophic levels, radically altering the global system’s overall equilibrium and resilience.
-($III_{A}$) System adaptation to multidirectional feedback loops.
-The third pathway feature is the global system adaptation to multidirectional feedback loops.
-This aspect focuses on the broad-scale adaptations of various subsystems to the impacts arising from multi-directional, diverse range of AIs.
-These adaptations occur through complex feedback loops that link AIs with multiple facets of broader societal, economic, and political subsystems.
-These loops entail reciprocal interactions, wherein AI-induced changes in one subsystem can both affect and be affected by shifts in other subsystems, resulting in a network of interconnected and mutually reinforcing effects.
-The accumulative AI x-risk manifests from the potential of these interlinked and reciprocal changes to progressively destabilize key subsystems.
-Although no single AI application may pose an immediate existential threat, the aggregated impact of AI across various adapted subsystems could lead to critical systemic imbalances or crises, when reaching a tipping point.
-[^14] One major challenge is to develop dynamic models that capture these shifting feedback mechanisms and to validate these models through both historical data and controlled experimentation.
-There remains uncertainty about how different types of feedback loops interact and how interventions can be designed to maintain a system within safe operational boundaries.
-Open questions pertain to the identification of early warning signals that might indicate a system is approaching a critical threshold where negative feedback can no longer contain an escalating process.
-Next steps include applying system dynamics to study the proposed concepts more concretely.
-The decisive ASI x-risk hypothesis is often illustrated using hypothetical scenarios like the well-known paperclip maximizer.
-To provide a more concrete and tangible representation of the accumulative AI x-risk hypothesis, I introduce a hypothetical scenario termed ”perfect storm MISTER.” [^15] The term ”perfect storm” refers to a situation where a rare combination of circumstances drastically aggravates an event.
+The next section applies this systems perspective to develop an illustrative case for the accumulative AI x-risk hypothesis: “The perfect storm MISTER” thought experiment.
+[^22] The term ”perfect storm” refers to a situation where a rare combination of circumstances drastically aggravates an event.
 It is typically used to describe scenarios in which a confluence of factors or events, which are individually manageable, come together to create an extraordinary and often catastrophic situation.
@@
 The story in both the book and film recounts a catastrophic weather event, where multiple meteorological elements converged to create a fierce and fatal storm.
-In broader applications, particularly in discussing complex systems or societal issues, ”perfect storm” characterizes situations where diverse negative factors or risks coalesce and interact.
+In broader applications, particularly in discussing systems or societal issues, ”perfect storm” characterizes situations where diverse negative factors or risks coalesce and interact.
 The interplay of these elements produces a compounded impact, far surpassing the severity one would expect from the sum of the individual parts.
@@
 The ”perfect storm” metaphor hence captures the essence of scenarios where the convergence of various challenges or risks creates a crisis of extraordinary magnitude.
-This scenario seeks to make the imaginative exercise of the accumulative hypothesis more tangible and relatable.
-In short, the perfect storm MISTER scenario encapsulates a progression of escalating social risks of Manipulation, Insecurity, Surveillance, Trust erosion, Economic destabilization, and Rights infringement.
-Each component in this scenario corresponds to a type of social problem and its associated AI social risk, as outlined in Section 2.
-The perfect storm MISTER scenario aims to illustrate how interlinked social risks can compound, surpassing critical thresholds, and consequently pose substantial existential threats.
+Each letter in MISTER represents a member of a subset of the social risks (Manipulation, Insecurity threats, Surveillance and erosion of Trust, Economic destabilization, and Rights infringement) that were introduced in Section 2.3.232323While not all social or ethical risks have existential implications, the Perfect Storm MISTER scenario focuses on those with potential for significant systemic impact.
+This will be followed by a comparison of decisive and accumulative AI x-risk pathways from a systems analysis perspective.
 4 The perfect storm MISTER
-Consider the highly interconnected world of 2035 where the pervasive integration of AI and IoT technologies has transformed almost every aspect of daily life.
-[^16] The Internet of Things (IoT) refers to the network of physical devices embedded with sensors, software, and other technologies for the purpose of connecting and exchanging data with other devices and systems over the Internet.
+Consider the highly interconnected world of 2040 where the pervasive integration of AI tools, AI assistants, AI agents, and Internet of Things (IoT) technologies has transformed almost every aspect of daily life.
+[^24] IoT refers to the network of physical devices embedded with sensors, software, and other technologies for the purpose of connecting and exchanging data with other devices and systems over the Internet (Rose et al., 2015; Li et al., 2015).
 These devices range from ordinary household items like refrigerators and thermostats to sophisticated industrial tools.
 Internet of things represents the idea of a highly interconnected world where real-time data exchange and automation are pervasive.
-Cities embody a higher level of automation as compared to today, various functions in domestic and industrial sectors are automated, and even the most mundane devices like mirrors and refrigerators have become part of a vast, data-exchanging network.
-Frontier AI systems have become the backbone of this connected world.
+Cities embody a higher level of automation as compared to today, with sustainability assistants (Rillig and Kasirzadeh, 2024) optimizing resource usage, AI agents (Xi et al., 2023) managing various functions in domestic and industrial sectors, and even the most mundane devices like mirrors and refrigerators have become part of a vast, data-exchanging network.
+Personalized AI assistants (Gabriel et al., 2024) have become the backbone of this connected world, serving roles from decision-making algorithms to social companions.
 However, beneath the surface of this technological connectivity, vulnerabilities and risks have been brewing.
-AI-driven manipulation.
-The capability of AI to create convincing deepfakes reaches a critical height, where distinguishing fact from deepfake-generated reality becomes nearly impossible.
-[^17] Concerns regarding AI-generated fake realities, especially in the context of manipulation and misinformation, have long been topics of research.
+Manipulation by AI assistants and agents.
+The abuse and misuse of AI systems for creating convincing deepfakes and misinformation reaches a critical point, where the information ecosystem becomes so polluted that rational public discourse becomes nearly impossible.
+[^25] Concerns regarding AI-generated fake realities, especially in the context of manipulation and misinformation, have long been topics of research (Whittaker et al., 2020; Sharma and Kaur, 2022).
 However, recent advancements in generative AI are introducing new dimensions to this issue.
@@
 Currently, no methods are entirely reliable in distinguishing these fabrications from actual reality.
-This capability opens the door to the creation of hyper-personalized, AI-generated propaganda and similar persuasive narratives, which can be strategically leveraged for social engineering purposes — manipulating public opinion, influencing electoral outcomes, inciting civil unrest, diminishing oppositional viewpoints, and exploiting existing social and epistemic divides.
-Proliferation of Insecurity threats.
-In the domestic sphere, everyday personal IoT devices like smart mirrors and refrigerators transcend their roles as mere conveniences.
-They now represent significant points of vulnerability.
-With frontier AI capabilities, cybercriminals can penetrate these devices in a more expansive way, leading to widespread identity theft and ushering in a new era of digital espionage.
-What were initially perceived as isolated incidents of breaches gradually coalesce into a discernible pattern, signifying a more amplified erosion of digital security.
-The expansion of IoT devices also paves the way for the creation of extensive, interconnected botnets.
-These AI-powered networks, once relatively benign, show agentic abilities and become capable of launching unprecedented Distributed Denial of Service attacks, against critical infrastructures, including national power grids and communication networks.
-They represent a significant leap in cybersecurity risks.
-Each attack, incrementally more sophisticated than the last, signifies a disturbing escalation from individual cybersecurity concerns to widespread threats against national security.
-Moreover, as AI technologies becomes more widely available, they facilitate the emergence of new forms of bioterrorism.
+The manipulation architecture operates through “cognitive cascade captures” (Hazrati and Ricci, 2024; Deldjoo et al., 2024) where initial successful manipulation creates vulnerabilities for subsequent influence attempts.
+Advanced AI technologies have enabled the creation of hyper-personalized propaganda and persuasive narratives, which can be strategically leveraged for social engineering purposes such as manipulating group identities, amplifying existing prejudices, exploiting belief systems through synthetic evidence, and creating personalized epistemic bubbles.
+Unlike traditional static influence attempts, AI assistants and agents embedded in recommender systems could maintain consistent manipulation pressure while dynamically adapting to individual and collective response patterns (Kasirzadeh and Evans, 2023; Carroll et al., 2024).
+This unprecedented technological capability, in turn, poses fundamental challenges for maintaining individual and collective autonomy, preserving shared reality frameworks, protecting democratic discourse, and ensuring societal resilience.
+Insecurity threats. The proliferation of IoT devices in domestic environments has fundamentally transformed the security landscape, creating unprecedented vulnerabilities in personal digital spaces.
+Digital-security threats.
+Smart devices, from mirrors to refrigerators, have evolved beyond their roles as mere conveniences to significant points of vulnerability.
+Cybercriminals can now penetrate these devices in increasingly sophisticated ways, leading to widespread identity theft and ushering in a new era of digital espionage.
+What were initially perceived as isolated breaches have gradually coalesced into a discernible pattern, signifying a more profound erosion of digital security.
+The expansion of IoT devices has simultaneously paved the way for the creation of extensive, interconnected botnets.
+These AI-powered networks, once relatively benign, now demonstrate agentic abilities and have become capable of launching unprecedented Distributed Denial of Service attacks against critical infrastructures, including national power grids and communication networks.
+Each attack, incrementally more sophisticated than the last, represents a disturbing escalation from individual cybersecurity concerns to widespread threats against national security.
+Bio-security threats.
+As AI technologies become more widely available, they facilitate the emergence of new forms of bioterrorism.
 Private research labs with minimal expertise in synthetic biology and chemistry are now using AI to develop more infectious and deadly pathogens.
@@
 Tools like retrosynthesis software, which assist in the design of molecules by reversing their synthetic processes, exemplify this trend.
-Global mass surveillance and erosion of trust.
-By 2035, the implications of AI-induced mass surveillance have profoundly transformed socio-political structures worldwide.
-The earlier instances, such as China’s social credit system and the controversial NSA’s mass data collection, were just the precursors to a now-global trend.
-Governments across various political spectra have embraced advanced AI for mass surveillance, frequently under the guise of national security.
-This landscape is marked by an array of pervasive surveillance systems, some overt and others clandestinely operated under national security pretexts.
-In this future context, the reach of government surveillance have gradually eroded the sanctity of privacy even in putative democracies.
-The constant monitoring and evaluation of citizens by AI systems instill a culture of self-censorship among citizens, diluting the diversity of thought and social trust crucial for healthy democracy.
-In several countries, this leads to a radical erosion of trust in media as well as public and private institutions.
+Epistemic insecurity.
+Advanced AI assistants and agents have introduced fundamental challenges to both public and private epistemic infrastructures.
+[^27] Following Milano and Prunkl (2024) and Kay et al.
+(2024), epistemic infrastructure includes the systems, institutions, and practices through which knowledge claims are verified, transmitted, and maintained within societies.
+In public domains, this has manifested through the systematic erosion of shared verification mechanisms, where traditional epistemic authorities face unprecedented challenges as synthetic content becomes increasingly indistinguishable from authentic documentation.
+This disruption strikes at the heart of societal knowledge validation systems, undermining established protocols for information verification and authentication.
+In the private sphere, personal communication channels—traditionally resistant to large-scale manipulation through social trust mechanisms—have become increasingly vulnerable to synthetic content that convincingly mimics trusted sources (Pennycook and Rand, 2021).
+This disruption extends beyond mere communication interference, affecting personal correspondence authenticity, private record verification, and the fundamental reliability of interpersonal trust mechanisms.
+The impact on private epistemic frameworks represents a significant shift in how individuals verify and validate information within their personal networks.
+Surveillance and erosion of Trust.
+The transformation of mass surveillance through AI represents one of the deepest shifts in the relationship between state power and individual liberty.
+What began as isolated initiatives have evolved into a global phenomenon that transcends traditional political classifications, fundamentally altering the perceptions of privacy, social cohesion, and democratic governance.
+The historical trajectory of mass surveillance has reached a troubling convergence between authoritarian and democratic governance models.
+Early warning signs emerged with China’s social credit system and the NSA’s PRISM program, but these now appear almost wide-spread.
+[^28] For an analysis of this evolution, see Lyon (2021) examination of how pandemic-era surveillance measures accelerated the adoption of AI-driven monitoring systems.
+Multiple earlier revelations have particularly warned against the erosion of democratic norms by AI-enabled surveillancce technologies such as spywares (Farrow, 2022; Rujevic, 2024).
+[^29] Farrow’s groundbreaking 2022 investigation exposed how putative democracies have embraced surveillance technologies traditionally associated with authoritarian regimes.
+The cases of Greece’s phone-hacking campaign targeting opposition politicians and Poland’s deployment of Pegasus spyware against civil society actors demonstrate how surveillance technologies can be weaponized even within democratic frameworks.
+The European Parliament’s 2024 Special Committee report on surveillance spyware documents numerous instances of democratic governments using these technologies against their own citizens.
+The societal implications of ubiquitous surveillance manifest in what Han (2015) describes as the “transparency society,” where the mere possibility of observation fundamentally alters social behavior.
+[^30] Han’s (2015) The Transparency Society provides a philosophical framework for understanding how constant surveillance reshapes social relations and individual psychology.
+Earlier studies (Kaminski and Witnov, 2014) investigated and document situations where citizens modify their behavior not in response to actual surveillance but to its perceived omnipresence.
+[^31] Murray et al.’s (2024) empirical study documents self-censorship across multiple societies, correlating directly with the implementation of AI surveillance systems.
+This chilling effect on public behavior and discourse represents a particularly insidious threat to democratic vitality, as it operates through self-imposed constraints rather than direct coercion.
 Economic destabilization.
-In the financial market infrastructure, AI algorithms increasingly play a substantial role in analyzing and simulating market trends.
-This progression, however, introduces new challenges to economic stability.
-One emerging concern is the use of AI in creating sophisticated, but deceptive financial products.
-These products, while not outright phantom instruments, are complex derivatives or investment vehicles.
-It becomes increasingly difficult to distinguish these products from more traditional, stable investments, leading to potential market disruptions.
-The impact of these AI-enhanced financial products is particularly significant in already vulnerable economies facing systemic issues and cyber threats.
-The introduction of these complex investment tools further undermines trust in financial systems, intensifying the risk of economic downturns.
-Additionally, the potential for AI to unintentionally trigger market instability becomes a realistic concern.
-Automated trading algorithms, responding rapidly to market signals, now can inunprecedented speed amplify market volatility at unprecedented speeds.
-This phenomenon is exemplified in incidents like flash crashes — i.e., sudden and drastic market fluctuations driven by the high-speed responses of automated trading systems.
+By 2040, the global economy has entered an unprecedented phase of instability, driven by the rapid and unmanaged deployment of AI systems across industries since the late 2020s.
+Nearly 40% of pre-2025 jobs have been eliminated by automation, creating massive structural unemployment.
+The pace and sophistication of AI adoption left no sufficient time for meaningful workforce transitions.
+The promised creation of new jobs never materialized at scale — AI systems became in charge of handling their own maintenance, optimization, and even creative development.
+The proposed solution of Universal Basic Income (UBI), while theoretically promising, has fallen short in practice due to political constraints and corporate resistance to the taxation necessary for meaningful implementation.
+[^32] See Auken (2016).
+The concentration of AI capabilities among a handful of tech conglomerates has exacerbated wealth inequality to historic levels, with the top 0.1% now controlling over 70% of global wealth.
+The economic upheaval has been amplified by the fragmentation of the global order into competing digital-industrial blocs.
+The US-led Western alliance and the China-centered Asian sphere have created parallel technological and financial ecosystems, effectively splitting the world economy.
+This digital iron curtain has disrupted decades of global trade integration, with companies forced to maintain separate systems and standards for each bloc.
+The emergence of competing digital currencies has undermined the dollar-based financial system, while AI-powered economic warfare — including automated sanctions enforcement, algorithmic trade restrictions, and digital blockades — has become a daily reality.
+Smaller nations find themselves forced to align with one bloc or risk economic isolation, further destabilizing regional powers and trade relationships.
+Market stability faces additional challenges from the acceleration of algorithmic trading systems.
+While high-frequency trading is not new, the integration of advanced AI capabilities introduces novel forms of systemic risk.
+These systems can now process and react to market signals at unprecedented speeds, potentially creating feedback loops that amplify market volatility.
+The phenomenon extends beyond simple flash crashes to what might be termed ”cascade failures,” where AI-driven trading systems interact in ways that create emergent instability patterns.
 Rights infringement.
-The pervasive application of AI in mass surveillance and extensive data collection practices is increasingly encroaching upon basic human rights.
+The pervasive application of AI in mass surveillance and extensive data collection practices has extensively encroached upon basic human rights.
 Privacy breaches have become alarmingly routine as AI systems gather and analyze personal data on an unprecedented scale.
@@
 Such practices, often lacking transparency and accountability, lead to systemic unjust treatment and exacerbate existing societal inequalities at scale.
-In the perfect storm MISTER Scenario, a series of interconnected AI-induced risks coalesce into a catastrophic sequence of events, each exacerbating the next, leading to a potential existential crisis for humanity.
+In the perfect storm MISTER Scenario, against this background condition, a series of interconnected AI-induced risks coalesce into a catastrophic sequence of events, each exacerbating the next, leading to an existential crisis for humanity.
 The AI x-catastrophe unfolds with a devastating AI-driven cyberattack simultaneously targeting critical power grids across three continents.
@@
 Nations or non-state actors, driven by desperation or opportunism, engage in aggressive military actions, potentially leveraging AI technologies in warfare without legal constraints.
-In this scenario, the x-catastrophe emerges from the synergistic failure of systems critical to the functioning and survival of human civilization.
+In this scenario, the x-catastrophe arise from the synergistic failure of systems critical to the functioning and survival of human civilization.
 The simultaneous and compounded nature of these crises creates a perfect storm situation where not only is recovery extremely challenging, but the potential for irreversible collapse is a stark reality.
-Before discussing the ramifications of the accumulative AI x-risk hypothesis, let’s first consider and address two potential objections.
+With the perfect storm MISTER scenario established, we now compare the causal pathways underlying decisive and accumulative hypotheses from a systems analysis perspective.
+5 Pathway to decisive ASI x-risk
+The decisive ASI x-risk hypothesis examines scenarios, like the paperclip maximizer, where ASI could abruptly trigger existential catastrophes.
+This hypothesis assumes catastrophic outcomes arise from a single, identifiable cause: the creation of a misaligned superintelligence capable of rapid system-wide disruption.
+First, ASI serves as the initial perturbation source, introducing novel disruptions into the global system.
+Second, the extensive connectivity of modern global infrastructures to ASI enables rapid propagation of ASI-initiated disruptions.
+Third, predominantly unidirectional dependencies prevent the system from self-correcting, instead reinforcing and accelerating the catastrophic trajectory.
+Figure 2: Pathway to decisive ASI x-catastrophe
+Figure 2 illustrates the causal pathway in decisive AI x-risk scenarios.
+Assumptions about the type of pathway for risk propagation are as follows.
+Modern civilization operates through densely connected subsystems — from financial markets and supply chains to communication infrastructures and social institutions.
+In the figure, square nodes represent various subsystems in the global world, with red squares indicating subsystems impacted by the ASI and grey squares showing those not yet impacted.
+Stronger impacts are shown with bold red links, while weaker or potential connections are indicated in grey.
+The arrangement follows a temporal sequence for illustrative purposes.
+In scenarios like the paperclip maximizer, what begins as a simple optimization goal (maximize paperclips) cascades through multiple subsystems as the ASI pursues instrumental subgoals like resource acquisition.
+Initial control of computational resources spreads through economic networks, infrastructure control cascades through technological systems, and responses to human resistance ripple through social and political networks.
+Unlike natural systems that have inherent balancing mechanisms (like predator populations limited by prey scarcity), an ASI creates self-reinforcing cycles without effective counterbalances.
+For example, the ASI’s acquisition of computing resources, as an instrumental subgoal, increases its capabilities, enabling more effective strategies for pursuing further subgoals.
+Each cycle amplifies the ASI’s ability to pursue its ultimate objective while simultaneously reducing the possibility of external intervention.
+With access to critically important components (from global manufacturing to energy grids), the ASI can leverage all necessary subsystems, with no significant subsystem immune from its influence (illustrated by the red diamond with x).
+This progression lacks natural checks - nothing inherently constrains its escalating influence as it pursues its final goal of converting resources into paperclips.
+6 Pathway to accumulative AI x-risk
+The accumulative AI x-risk hypothesis posits that existential risks may arise from multiple interacting disruptions that compound over time, progressively weakening systemic resilience until a triggering event causes unrecoverable collapse.
+Unlike decisive scenarios with a single causal path, this hypothesis involves multiple causal processes that collectively contribute to existential risk.
+First, different types of AI systems serve as initial perturbation sources, where localized impacts, even if minor at the outset, can aggregate and intensify across various subsystems.
+Second, while modern infrastructures are highly interconnected, AI systems typically impact specific subsystem clusters — creating selective pathways for disruptions to propagate through the network rather than the pervasive reach by ASI seen in decisive AI x-risk scenarios.
+Third, disruptions from different AI clusters interact and amplify each other as they spread through connected subsystems, creating cumulative effects larger than their initial impacts.
+Finally, as impacts from multiple AI clusters accumulate across subsystems, the system’s capacity for self-correction diminishes, making each new disruption more likely to reinforce rather than resolve existing instabilities.
+Figure 3: Pathway to accumulative AI x-catastrophe
+Figure 3 illustrates the causal pathway in accumulative AI x-risk scenarios.
+Each hexagon represents a cluster of problems compounded by AI assistants or agents in perfect storm MISTER: Manipulation by AI assistants and agents, Insecurity threats, Surveillance and erosion of Trust, Economic destabilization, and Rights infringement.
+Within the hexagons, circles and squares represent different entities (institutions, humans, communities) interacting with AI systems.
+The accumulative AI x-risk manifests itself from the potential of these interlinked and reciprocal changes to progressively destabilize key subsystems.
+Although no single AI application poses an immediate existential threat, the aggregated impact of AI across various adapted subsystems leads to critical systemic imbalances or crises (diamonds in red).
+[^33] One major challenge is to develop dynamic models that capture these shifting feedback mechanisms and to validate these models through both historical data and controlled experimentation.
+There remains uncertainty about how different types of feedback loops interact and how interventions can be designed to maintain a system within safe operational boundaries.
+Open questions pertain to the identification of early warning signals that might indicate a system is approaching a critical threshold where negative feedback can no longer contain an escalating process.
+Next steps include applying system dynamics to study the proposed concepts more concretely.
+To summarize, the decisive and accumulative AI x-risk hypotheses differ in three key aspects.
+First, in perturbation source: a single ASI versus multiple AI systems creating localized disruptions.
+Second, in propagation pattern: the ASI’s pervasive reach enabling rapid system-wide cascades versus selective pathways through specific subsystem clusters.
+Third, in catastrophic development: immediate unidirectional acceleration versus gradual accumulation of interacting disruptions that progressively degrade system resilience.
+While both pathways can lead to system-wide catastrophe, they do so through fundamentally different causal mechanisms.
+Before examining the implications of the accumulative AI x-risk hypothesis for governance and long-term safety, let us address several potential objections.
 7 Objections and replies
-Objection 1: The societal breakdown or accumulative civilizational collapse, as it is characterized, does not equate to the x-catastrophes often envisioned in AI risk discussions, such as an ASI power-seeking entity deliberately aims to destruct humanity.
-Historically, civilizational collapses have happened and, while significant, have not equated to human extinction or unrecoverable collapse.
+Objection 1: The societal breakdown or accumulative civilizational collapse, as characterized here, does not equate to the x-catastrophes often envisioned in AI risk discussions, such as an ASI power-seeking entity deliberately aiming to destroy humanity.
+Historically, civilizational collapses have occurred and, although significant, have not been equated to human extinction or unrecoverable collapse.
 Reply 1: This objection, while historically grounded, may not fully consider the unique AI x-catastrophic threats posed in our modern, interconnected global civilization.
 The modern world functions as a tightly interconnected global system, far more integrated than any past civilization.
-The COVID-19 pandemic serves as a contemporary example, where a health crisis originating in one region quickly escalated into a global emergency, disrupting economies, supply chains, and everyday life worldwide.
+The COVID-19 pandemic serves as a contemporary example, where a health crisis originating in one region quickly escalated into a global emergency, disrupting economies, supply chains, and everyday life around the world.
 The perfect storm MISTER example illustrates how a collapse in one sector or region can have rapid, global impacts in today’s interconnected world, a phenomenon not seen in past civilizational collapses.
@@
 Reply 2: The complexity inherent in the accumulative AI x-risk model is actually a key advantage.
-It promotes a comprehensive understanding of the diverse impacts of AI across societal, economic, and ecological domains.
-By acknowledging the interdependencies and potential for cascading effects within these systems, we can increase resilience and adaptability against x-catastrophic disruptions, leading to the development of more robust long-term strategies.
+By acknowledging the interdependencies and potential for cascading effects across societal, economic, and ecological domains, we can develop more targeted monitoring mechanisms.
+These mechanisms would track critical thresholds, measure compounding effects, and identify early warning signals of system destabilization.
+While such monitoring systems require significant development, they are essential for understanding how AI impacts accumulate across complex global systems.
 While the paperclip maximizer example offers a simple and elegant model, it deeply overlooks the nuanced realities of AI’s interactions with complex global systems.
 The accumulative model, in contrast, acknowledges this complexity and provides a more realistic portrayal of potential risks.
-It emphasizes detailed, empirically-grounded analysis, which is crucial for identifying critical areas sensitive to existential risks.
-Future efforts should aim to further validate and refine the accumulative hypothesis through system dynamic simulations (e.g., Karnopp et al.
+It calls for detailed, empirically-grounded monitoring and analysis, which is crucial for identifying critical areas sensitive to existential risks.
+In addition to establishing monitoring mechanisms, future efforts should aim to further validate and refine the accumulative hypothesis through system dynamic simulations (e.g., Karnopp et al.
 (2012)).
-These simulations would offer tangible insights into the complex interplay of AI-induced risks, enhancing our understanding and ability to model these interactions accurately.
-Although this paper does not resolve all the questions surrounding the accumulative model, it shows the need for ongoing research and deeper exploration in the effective management of AI x-risks.
-The approach towards a more nuanced understanding of AI’s potential impact on our world is both necessary and urgent.
-8 Implications
-In this section, I briefly discuss five governance implications of this expanded conception of AI x-risks.
-8.1 Consolidated approach to AI risk governance
-Recognizing the distinct nature of decisive and accumulative AI x-risks calls for a holistic evaluation of how we govern AI risks.
-The governance of decisive ASI x-risks may be best served by centralized and stringent control measures.
-These controls aim to quickly address and mitigate the immediate and widespread consequences that such particular risks could unleash.
-For instance, stringent regulation of ASI research and development is needed.
-This might include international agreements akin to the Nuclear Non-Proliferation Treaty, ensuring that advancements towards ASI are closely monitored and controlled.
-Rapid-response protocols, similar to those used in handling biochemical threats, could be implemented.
-In the event of an ASI-related crisis, a centralized authority could quickly mobilize resources and expertise to mitigate the threat.
-On the other hand, accumulative AI x-risks, which develop gradually and are often less perceptible in the short term, calls for decentralized, adaptable governance.
-This could involve a network of local and regional AI oversight bodies that monitor the impacts of AI in their respective areas, akin to environmental protection agencies that oversee local environmental concerns.
-The unification of different domains in AI risk governance not only streamlines efforts but also avoids the unnecessary costs associated with rebranding.
-AI development is a gradual and continuous process, and similarly, the development of tools for detecting and mitigating AI risks should also be progressive and sustained.
-The integrated approach prevents the redundant effort of ’reinventing the wheel’ for each distinct aspect of AI x-risk mitigation.
-To clarify, the consolidated approach to AI risk governance does not necessitate a direct link between every project and each facet of social risks to x-risks.
-Rather, it advocates for a meta-view or a macro-level perspective.
-This perspective facilitates a comprehensive understanding of how various projects, while not necessarily directly interlinked, collectively contribute to the overarching objective of mitigating AI x-risks.
-In light of this discussion, I propose to establish a dynamic ”risk contribution checklist.” This checklist, designed to be both flexible and regularly updated, would serve as a tool for researchers and institutions engaged in AI development.
-It would enable a coordinated approach to risk mitigation, ensuring that diverse projects align with and contribute effectively to the broader goal of managing AI existential risks.
-8.2 Harmonization of ethical and existential risks
-Traditionally, particularly in the Anglo-American context, there has been a tendency to treat communities focusing on ethical risks and those concerned with x-risks as distinct (Section 2).
-However, the accumulative AI x-risk hypotheses challenge this separation, indicating that such a dichotomy is epistemically unsound in the context of AI risks.
-In addition, the distinction between decisive and accumulative AI x-risks suggests a potential realignment of AI x-safety research priorities.
-While investigating catastrophic failure modes associated with ASI remains crucial, there is a growing need to allocate comparable attention and resources to the exploration of more subtle, long-term existential effects of AI via ethical risks.
-These effects, often cumulative in nature, require thorough study to understand their full impact and implications.
-Neglecting the accumulative aspect of AI x-risks would be a critical oversight.
-Failing to recognize and investigate these gradual and accumulative impacts could leave us unprepared for the nuanced existential challenges.
-Moreover, the harmonization facilitates the transfer of risk mitigation strategies across different projects.
-A prime example of this is the evolution of interpretable models to mitigate social AI risks into mechanistic interpretability to mitigate AI x-risks.
-This progression demonstrates how methodologies initially developed in one domain can be adapted and enhanced for application in another, thereby ensuring a more cohesive and efficient approach to managing AI risks.
-8.3 Short-long-term risk management
-In the literature, there has been an implicit assumption that it is plausible to separate short-term (ethical) and long-term (existential) AI risks.
-As I have suggested, upon closer examination, these lines blur, revealing a complex interplay of concerns that transcend traditional time frames.
-The immediate risks of advanced AI systems is not solely confined to the present but also encompasses the rapidly approaching future, indicating the interconnected nature of these risks.
-As such, the structure of AI risk-related problems indeed represents a complex blend of immediate, intermediate, and long-term considerations, challenging our conventional understanding of risk and time.
-Recognising the accumulative emergence of x-risks as a result of mismanagement, accumulation, or contingency path through ethical risks allows us to plan for schedules that address simultaneously both short-term and long-term risk mitigations and emphasise the importance of ongoing monitoring and evaluation.
+These simulations would offer tangible insights into the complex interplay of AI-induced risks.
+Although this paper does not resolve all the questions surrounding the accumulative model, it shows the need for ongoing research and deeper exploration in the effective conceptualization and management of AI x-risks.
+8 Risk governance and accumulative risk hypothesis
+The imperative to govern AI has become almost synonymous with mitigating its risks (Kaminski, 2023).
+The risk-based approach to AI governance is evident in major policy frameworks such as the NIST AI Risk Management Framework (National Institute of Standards and Technology, 2024), which presents risk taxonomies that align with earlier classifications (Bender, 2013; Weidinger et al., 2021; Bird et al., 2023) and regulatory approaches like the EU AI Act’s risk-based categorization.
+But how do different conceptions of AI risk fit together in risk governance efforts?
+The risk landscape is fragmented across multiple taxonomies with calls for governing social and ethical risks, catastrophic risks (Kasirzadeh, 2024), extreme risks (Shevlane et al., 2023; Bengio et al., 2024), and existential risks.
+These separate terminologies create what we might term “the risk fragmentation problem” — where distinct approaches to conceptualizing AI risk fail to provide complete coverage, leaving dangerous blind spots where risks can accumulate unnoticed.
+This fragmentation manifests in several ways.
+Social and ethical risk frameworks operate in isolation from frameworks addressing catastrophic risks (Anthropic, 2023).
+Immediate risk assessments rarely connect with analyses of long-term societal implications.
+Assessment methodologies remain siloed within their specific domains and terminologies, creating spaces where risks go unmonitored or unaddressed.
+[^36] There are at least four principal models for risk governance, each representing different balances between quantitative rigor and democratic accountability (Kaminski, 2023).
+First, the U.S.
+administrative model (1960s-1980s) emphasizes heavily quantitative risk assessment methodologies, prioritizing cost-benefit analysis and measurable outcomes (Boyd, 2012).
+This approach, while offering analytical precision, often struggles to account for uncertainties and non-quantifiable or hardly-quantifiable risks that characterize emerging new technologies.
+[^34] This tension between quantification and uncertainty becomes particularly acute when dealing with novel AI risks that lack historical precedent for statistical analysis.
+Second, the democratic oversight model, exemplified by the National Environmental Policy Act (NEPA), emphasizes public participation and transparent decision-making processes (Froomkin, 2015; Kaminski and Malgieri, 2020).
+This approach can incorporate precautionary principles, acknowledging that when facing potentially catastrophic risks, the absence of complete scientific certainty should not preclude protective measures.
+[^35] The precautionary approach becomes especially relevant for AI governance given the potential for irreversible societal impacts.
+Third, the centralized risk evaluator assesses risk on a macro-level (Hampton, 2005; Black and Baldwin, 2010).
+Regulators identify the risk to be managed, select a level of risk tolerance, assess the harms and the likelihood of their occurrence, assign risk scores to firms or activities (such as “high,” “medium,” or “low”), and link the allocation of enforcement and inspection resources to risk scores.
+The Draft EU AI Act is a clear descendant of this kind of law.
+Fourth, enterprise risk management approaches, as represented by NIST standards (National Institute of Standards and Technology, 2024) and frontier model safety frameworks (Kasirzadeh, 2024), focus on how companies can organize internally to mitigate their risks.
+They may conduct ongoing risk analysis and mitigation to avoid liability or other penalties, whether regulatory or market-based.
+While enterprise risk management can occur in the absence or shadow of law, regulators can also participate by nudging companies to conduct risk mitigation through oversight, through the threat of regulatory enforcement, by offering safe harbors, or by issuing best practices or other guidance.
+Enterprise risk management is typically (1) cyclical and ongoing, and (2) organizational in nature.
+These models diverge not only in their approaches to hard versus soft law but also in their mechanisms for ensuring accountability and their treatment of uncertainty.
+While quantitative cost-benefit analysis often dominates traditional risk assessment, the unique challenges posed by AI technologies suggest the need for hybrid approaches that can incorporate both precautionary principles and rigorous analytical methods.
+Recent developments in AI risk assessment suggest a growing recognition of the need to combine multiple governance approaches to address the full spectrum of potential risks.
+The accumulative conception of AI x-risk — or its weaker interpretation as accumulative AI catastrophic risk — provides a unifying framework that can help bridge fragmented risk governance approaches.
+By emphasizing how risks compound and interact across domains, this perspective incentivizes connecting previously siloed risk frameworks.
+8.1 Holistic approach to AI x-risk governance
+The distinction between decisive and accumulative AI x-risks requires different, complementary, governance approaches.
+Decisive ASI x-risks call for centralized control measures similar to nuclear non-proliferation frameworks.
+This includes international monitoring of advanced AI development, strict development protocols, and coordinated emergency response mechanisms.
+The potential for rapid, system-wide impacts necessitates unified oversight and quick response capabilities.
+Accumulative x-risks, by contrast, require distributed monitoring systems that can track how multiple AI impacts compound across different domains.
+This suggests a network of oversight bodies monitoring specific sectors and subsystems, while sharing data about emerging risk patterns.
+Like financial regulators tracking systemic risk, these bodies would need mechanisms to detect when accumulated disruptions approach critical thresholds.
+These different governance needs can be integrated through a tiered framework: distributed monitoring for accumulative risks, coupled with centralized oversight for advanced AI development.
+This approach leverages existing governance structures while adding new capabilities for tracking risk accumulation.
+8.2 Unifying social and existential risks
+Traditionally, there has been a tendency to treat concerns focused on ethical risks and those concerned with x-risks as distinct (see Section 2).
+However, the accumulative AI x-risk hypotheses challenge this assumed separation, indicating that such a dichotomy is epistemically unsound in the context of AI x-risks.
+The distinction between decisive and accumulative AI x-risks suggests a needed rebalancing of safety research priorities.
+While investigating ASI failure modes remains important, equal attention should be given to understanding how social risks compound into existential threats.
+These compounding effects demand systematic study to understand how they might escalate into existential threats.
+Focusing solely on decisive scenarios while neglecting accumulative pathways would leave us blind to critical risks that build gradually through system interactions.
+Risk framework unification enables critical risk mitigation methodologies to transfer between domains.
+The evolution from interpretable models for social risks to mechanistic interpretability for existential risks demonstrates this potential.
+However, we must systematically identify what additional components are needed to scale current ethical frameworks to address catastrophic risks.
+Recognizing the accumulative emergence of x-risks as a result of mismanagement, accumulation, or contingency path through ethical risks allows us to plan for schedules that address simultaneously both short-term and long-term risk mitigations and emphasise the importance of ongoing monitoring and evaluation.
 It is interesting to explore how current AI risk management frameworks (Baryannis et al., 2019; Tabassi, 2023) adapt to the accumulative and decisive hypothesis in this paper.
-8.4 Enhancing Understanding and Awareness of AI Risks
-The critical link between non-existential social risks and x-risks shows the importance of elevating public awareness and education regarding the radical impacts of AI.
-This effort is geared towards empowering individuals and communities to participate meaningfully in conversations surrounding responsible AI development.
-Decisive x-risks, characterized by their immediate and striking effects, has tended to capture public attention.
-They resonate with cognitive biases, tap into inherent fears, or are amplified by sensationalist media portrayals.
-This dynamic may result in a distorted perception of AI x-risks, where the immediacy and drama of these risks eclipse the more subtle and gradually developing accumulative x-risks.
-In contrast, the accumulative risks, owing to their incremental nature and less obvious long-term impacts, often do not generate the same level of public interest or urgency.
-They typically do not provoke the instant emotional reaction seen with decisive x-risks, potentially leading to delayed public recognition and response until these risks escalate to critical levels.
-Therefore, it is important to steer the conversation on AI risks away from purely dystopian or utopian narratives.
-While such narratives may be engaging, they frequently lack the necessary depth and complexity for constructive discussion and effective risk management.
-A more nuanced and informed approach to discussing AI risks is crucial for fostering a balanced understanding and preparing for the future with AI.
-8.5 Mitigating x-catastrophes by limited connectivity
-Finally, the distinction between two types of AI x-risks explores the strategy of mitigating decisive AI-related x-catastrophes by deliberately avoiding universal connectivity to ASI.
-Understanding the risk landscapes through systems analysis reveals the potential benefits of establishing multiple, critical nodes.
-This strategic decentralization serves as a protective measure, enhancing global resilience against the spread of catastrophic events, particularly those instigated by ASI.
-If the paper clip maximizer is not expansively connected to humanity, the potential for decisive AI x-risk substantially diminishes.
-Consequently, the discussion so far advocates for a cautious stance against the globalization of ASI.
 9 Concluding remarks
-I began by asking whether the conception of x-risk from AI systems relate to the conception of non-existential social risks.
-Conventional discussions of AI x-risks often respond negatively to this question, and overlook an alternative way by which AI could bring about x-catastrophes.
-Against this background, I formulated an alternative view of AI x-risk.
-I used systems analysis perspective to differentiate between the two conceptions of AI x-risks in relation to two distinct causal paths that could have AI x-catastrophes as their endpoint.
-The two type of causal paths differ based on the way they instantiate three path features: non-linearity, connectedness, and dynamic system adaptation.
-Causal pathways that feature humanity cascade effect of ASI, super-connectedness of infrastructures to ASI, and system adaptation to unidirectional feedback loops, can end in decisive ASI x-catastrophes.
-Causal pathways that feature local causality with compounded systemic impact, selective-connectedness of infrastructures, and dynamic system adaptation to multidirectional feedback loops, can end in accumulative AI x-catastrophes.
-So x-risks should be about those events that occur in pathways and mitigation strategies targetted at them.
-I used a scenario termed perfect storm MISTER to propose how the accumulative model can give rise to AI x-catastrophes.
-I also discussed five governance implications of my proposed expanded conception of AI x-risks.
-Going forward, I think there is no inherent reason to consider that the accumulative hypothesis is any less likely than the decisive view.
-The need to further substantiate the accumulative hypothesis is apparent.
-This could be effectively addressed through the development of computational simulations using a system dynamics approach, aligning with the perspective I have established in analyzing the two types of AI existential risks.
-Finally, my discussion has so far only differentiated between two types of AI x-risks.
-The possibility of additional categories remains an open question.
-The implications of my arguments for assessing the probabilities of catastrophic outcomes and existential risks are significant and warrant further exploration in future research.
+This paper critically examined how AI existential risks are conceptualized in the current literature.
+The dominant framing, which I term the decisive AI x-risk hypothesis, focuses on scenarios where AGI or ASI directly causes catastrophic outcomes at a decisive moment, exemplified by thought experiments like the paperclip maximizer.
+Through systems analysis, I proposed and defended an alternative causal pathway to AI existential catastrophes: the accumulative AI x-risk hypothesis.
+This hypothesis suggests that existential risks can emerge from the compounding of multiple lower-severity disruptions that gradually weaken systemic resilience until a triggering event causes unrecoverable collapse.
+The perfect storm MISTER scenario illustrates this pathway, demonstrating how interactions between different types of AI risks could lead to existential catastrophe through progressive system degradation rather than just a single decisive event.
+This accumulative perspective has important implications for AI risk governance.
+While the decisive view tends to treat x-risks as categorically distinct from other AI risks, the accumulative hypothesis reveals critical relationships between different risk types and their potential to compound.
+This suggests the need for integrated governance approaches that address both immediate AI risks and their potential accumulative effects.
+The systems analysis approach shows how seemingly manageable risks can interact and amplify through feedback loops and network effects, creating emergent threats to systemic stability that may be overlooked when risks are analyzed in isolation.
+Several key questions remain from this analysis that warrant further investigation.
+First, we need better methods for identifying when disruptions become critically significant.
+This requires developing sophisticated monitoring systems that can track not just individual AI incidents, but also their cumulative effects and interactions across different subsystems.
+Such monitoring systems would need to establish clear thresholds that signal when accumulating disruptions are approaching dangerous levels.
+Second, while this paper introduces systems analysis to AI x-risk conceptualization and evaluation, we need more structured approaches for analyzing how risks accumulate.
+This includes developing formal frameworks for mapping causal chains, identifying key feedback loops, and understanding how different types of AI risks interact within complex global systems.
+Such frameworks would help systematically identify potential accumulative pathways to existential risk that might be overlooked by current approaches.
+Third, the quantification of accumulative AI x-risks presents unique methodological challenges.
+Unlike decisive scenarios where single events trigger catastrophic outcomes, accumulative risks involve complex interactions over time that are harder to model mathematically.
+We need new methods for calculating how multiple smaller risks combine and amplify, and how to assess the probability of system-wide failures emerging from these interactions.
+These challenges require innovations in risk modeling that can capture the dynamic, non-linear nature of risk accumulation.
+Looking ahead, there is no inherent reason to consider the accumulative hypothesis less plausible than the decisive view.
+Future work should focus on developing computational simulations using system dynamics to further substantiate the accumulative hypothesis and explore its implications.
+While this paper has contrasted decisive and accumulative pathways, other potential causal pathways to AI x-risk may exist.
+The framework developed here provides a foundation for investigating these possibilities through formal modeling of how different AI risks might interact and compound over time.
+Acknowledgments.
+I thank Mazviita Chirimuuta, Iason Gabriel, Jan Matusiewicz, Mario Guenther, Matthijs Maas, and John Zerilli for their valuable feedback.
+I am also grateful to participants in the AI, Data & Society group at the University of Edinburgh, the Ethics Reading Group at the University of Edinburgh, the Workshop on Digital Democracy at the University of Zurich, the Sociotechnical AI Safety Workshop in Rio de Janeiro, and members of the Collective Intelligence Institute.
+I thank Bálint Gyevnár for helping to create the figures.